### v1

# PyLO: Towards Accessible Learned Optimizers in PyTorch

Abstract
Learned optimizers have been an active research topic over the past decade, with increasing progress toward practical, general-purpose optimizers that can serve as drop-in replacements for widely used methods like Adam. However, recent advances—such as VeLO, which was meta-trained for 4000 TPU-months, remain largely inaccessible to the broader community, in part due to their reliance on JAX and the absence of user-friendly packages for applying the optimizers after meta-training. To address this gap, we introduce PyLO, a PyTorch-based library that brings learned optimizers to the broader machine learning community through familiar, widely adopted workflows. Unlike prior work focused on synthetic or convex tasks, our emphasis is on applying learned optimization to real-world large-scale pre-training tasks. Our release includes a CUDA-accelerated version of the small_fc_lopt learned optimizer architecture from (Metz et al., 2022a), delivering substantial speedups—from $39.36$ to $205.59$ samples/sec throughput for training ViT-B/16 with batch size $32$. PyLO also allows us to easily combine learned optimizers with existing optimization tools such as learning rate schedules and weight decay. When doing so, we find that learned optimizers can substantially benefit. Our code is available at https://github.com/Belilovsky-Lab/pylo
Machine Learning, ICML
Figure 1: Training step timing breakdown for ViT-B/16 on a single A100GPU. We measure the forward (blue), backward (orange), and optimizer (other colors) times per training step. We observe that the CUDA-accelerated learned optimizer step shows substantial improvements over the naive implementation (red) and is competitive with adafactor (purple). In all cases, as the batch size is increased, the relative overhead of the optimizer shrinks.
1 Introduction
Figure 2: PyLO: simplifies the integration of learned optimizers into standard machine learning workflows. By addressing key usability challenges, PyLO provides seamless access to meta-learned optimization techniques through three core features: (1) automatic weight loading from Hugging Face Hub, (2) a familiar PyTorch-style optimizer interface, and (3) accelerated CUDA kernel support. The library bridges the gap between advanced meta-learning research and practical machine-learning applications, enabling researchers and practitioners to easily leverage state-of-the-art learned optimization techniques in Pytorch.
Learned optimization (LO) (Hochreiter et al., 2001; Andrychowicz et al., 2016) represents a promising yet underutilized direction for advancing optimization algorithms in machine learning. Training contemporary neural networks requires solving highly non-convex optimization problems for which theory neither guarantees convergence to a globally optimal solution, nor convergence at the optimal rate. Therefore, despite the purportedly strong performance of hand-designed optimizers such as Adam (Kingma & Ba, 2017), for many tasks of interest, there may be substantial room for improvement. Therefore, data-driven optimization algorithms have the potential to make substantial improvements. However, even the most performant publically available learned optimizer to date, VeLO (Metz et al., 2022b), has seldom been adopted by deep learning practitioners.
Several critical barriers have hindered the more widespread adoption of learned optimizers: 1) an overfocus on meta-learning in current learned optimization frameworks, 2) the absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, 4) the absence of learned optimizer benchmarking for popular machine learning tasks, and 5) learned optimizer step overhead.
In what follows, we take a step towards improving the adoption of learned optimizers for deep learning by providing a PyTorch implementation of state-of-the-art learned optimizers that emphasizes their performance and accessibility for practical machine learning problems. our contributions can be summarized as follows:
  1. 1.
    We provide a modular open-source implementation of state-of-the-art learned optimizers in PyTorch which seamlessly integrates with torch.optim.Optimizer and the Huggingface ecosystem to enable easy usage within existing code and standardized sharing of task-specific learned optimizer weights.
  2. 2.
    We provide a CUDA-accelerated implementation of small_fc_lopt’s optimizer step which substantially reduces memory overhead and increases occupancy (GPU utilization as in Figure 1).
  3. 3.
    We benchmark optimizer step times and performance in PyLO for popular image classification and language modeling workloads, showing that our implementation scales to large workloads and improves on the existing JAX implementation.
  4. 4.
    Our flexible framework makes it easy to study the effect of weight decay and learning rate schedules on learned optimizer performance and, when doing so, we find that making these simple additions can substantially improve performance in some cases.
2 Project vision
With our goal of improving the adoption of learned optimizers within the deep learning community and enhancing the open-source ecosystem around them, we design PyLO with the following principles in mind:
Accessibility With minimal dependencies (PyTorch + HuggingFace) and a straightforward installation process, PyLO eliminates the technical barriers that have impeded the widespread adoption of learned optimizers thus far.
Decoupling By exclusively providing efficient learned optimizer implementations in PyTorch without meta-training code, PyLO decouples meta-testing from meta-training. This substantially reduces code bloat and additional dependencies required compared to other frameworks. As a result, PyLO is simpler to use and easier to understand.
Interoperability Through strict adherence to the PyTorch optimizer (torch.optim.Optimizer) interface specification, PyLO ensures seamless integration with existing training pipelines, enabling practitioners to adopt learned optimizers with minimal modification to existing code. It also helps to use external learning rate schedulers and decoupled weight decay.
3 Code design
We implement a modular architecture partitioned into three main components to realize our project vision. This design philosophy prioritizes both usability for practitioners and extensibility for researchers developing novel learned optimizers.
Optimization Module (pylo.optim) This module facilitates critical state management functions necessary for learned optimization. These include maintaining parameter-specific accumulators, computing parameter features for optimizer forward passes, executing update steps with configurable learning rate scaling, and supporting standard PyTorch optimizer functionality such as state dictionaries for resuming. This implementation allows practitioners to leverage advanced learned optimization techniques without significant modifications to existing training setups.
Meta-Model Architectures (pylo.models) encapsulates the parameters of the learned optimizer and its forward pass. It is also fully integrated with HuggingFaceHub (HuggingFace, ), allowing users to download existing optimizers from the hub and providing a mechanism for weight distribution and versioning of future optimizers. Through HF integration, we facilitate community-driven improvement cycles and allow researchers to easily leverage our CUDA-accelerated implementation for their own benchmarking.
CUDA Acceleration (pylo.csrc) The unique computational demands of learned optimizers present challenges that default PyTorch is unequipped to handle. Therefore, we implemented a specialized CUDA kernel for the small_fc_lopt learned optimizer. This implementation strategically leverages the GPU memory hierarchy to address the primary bottleneck: memory bandwidth rather than computational throughput. Our CUDA implementation employs two key optimization strategies: (1) Register-Based Computation and (2) On-the-Fly Feature Computation. The first utilizes GPU registers to store intermediate activations during forward propagation through the optimizer’s MLP. This approach minimizes high-latency global memory accesses by keeping frequently accessed data in the fastest available memory tier. The second avoids storing normalized features in global memory, recalculating them when needed, and trading redundant computation for reduced memory bandwidth.
Figure 3: Learned Optimizer step time scaling plot. We compare the optimizer step time for CUDA and JAX implementations of small_fc_lopt. We observe that our CUDA implementation allows scaling to larger models and substantially reduces step time.
4 Benchmarking LO step times in PyLO
In this section, we establish the per-step overhead of learned optimizers for common pre-training workloads and showcase the improved scaling behavior of PyLO’s CUDA-accelerated learned optimizer steps. Our goal is to report learned optimizer overhead relative to hand-designed optimizers for practical tasks and to illustrate how the overhead scales as the model size is increased. To this end, we benchmark and record the forward, backward, and optimizer step times for training ViT-B/16 and a $410$M parameter language model on a single $80$GB A100 GPU.
Learned Optimizer overhead shrinks as batch size increases Figure 1 reports timings at different batch sizes for ViT-B/16. We observe that as batch size is increased the overhead of the learned optimizer’s step compared to the forward and backward pass shrinks. This suggests that large-scale training with long per-step times may be a suitable workload for learned optimizers, where the cost of applying the optimizer can be amortized. Table 1 compares timing results across different optimizers for ViT and GPT. We observe that the per-step time small_fc_lopt (CUDA) increases due to the large parameter count of the language model but that using our CUDA kernels still leads to a remarkable improvement over the naive implementation. We note that it also halves the optimizer step time of the original JAX implementation.
Scaling behavior of our CUDA implementation Figure 3 reports the scaling behavior w.r.t. transformer size of different small_fc_lopt (we use hidden size $32$ as in (Thérien et al., 2024)) implementations: CUDA and JAX. JAX corresponds to the original implementation of (Metz et al., 2022a) which uses the JAX compiler, while CUDA is the fastest implementation that we make available in PyLO. We observe that our CUDA implementation results in a substantial memory savings as it can optimize the large model size while the JAX implementation fail for the 1B parameter transformer. Moreover, our CUDA implementation also outperforms the JAX compiler by 2x, which corresponds to a substantial reduction in optimizer overhead.
Model Optimizer BS/SL Fwd (ms) Bwd (ms) Opt step (ms) #Params (M)
ViT-B/16 AdamW 32/197 17.52 38.13 4.90 86.57
Adafactor 18.99
small_fc_lopt (naive) 756.80
small_fc_lopt (CUDA) 99.59
GPT2-355M AdamW 4/1024 197.41 392.29 20.12 355.92
Adafactor 35.11
small_fc_lopt (naive) 2872.17
small_fc_lopt (CUDA) 319.14
Table 1: Timing results across different optimizers for ViT and GPT. We report timings for realistic batch size (BS) and sequence length (SL) which fit on a single A100 GPU.
5 Demonstrating the real world use of PyLO
In the following section, we evaluate and compare our PyTorch implementation of VeLO (Metz et al., 2022b) and $\mu$LOM (Thérien et al., 2024) to the performance of tuned AdamW with a cosine annealing schedule. Our goal is to establish the performance of readily available learned optimizers in PyLO for these practical vision and language pre-training tasks.
(a) LM $\mu$LO
(b) LM VeLO
(c) ViT $\mu$LO
(d) ViT VeLO
Figure 4: Weight decay and learning rate schedule (LRS) ablation. In PyLO, it is easy to equip VeLO and $\mu$LOM with weight decay and learning rate schedules. We sweep different values and find that in some cases learned optimizers improve.
5.1 Training Vision Transformer on ImageNet-1K
Experimental Configuration: We train ViT-Base/16 models (86M parameters) for 480 epochs (150k steps) on ImageNet-1K following established protocols (Wightman, 2019). We apply standard augmentation techniques (RandAugment, CutMix, Mixup) to ensure realistic training conditions that reflect practical deployment scenarios. We employ a batch size of 4096 and conduct these experiments on Nvidia H100 GPUs.
Results As shown in Table 2 VeLO demonstrates competitive performance in this context, achieving 78.45% top-1 accuracy compared to AdamW’s 77.22%. This scenario demonstrates a case where the learned optimizer outperforms Adam. $\mu$LO exhibits training divergence after 65,000 steps due to its limited 5,000-step meta-training horizon, achieving only 62.14% peak validation accuracy.
Table 2: Final Loss for Language Modeling (LM) and Final Validation Accuracy for Vision Transformer (ViT) Tasks for Medium Model Sizes Using Different Optimizers
[…]
$\mu$LOM 3.18 62.14
VeLO 2.89 78.39
AdamW + Cosine 2.91 77.22
5.2 Language model pre-training on FineWeb
Experimental Configuration We pre-train a $355$M parameter GPT-2 style transformers on FineWeb-EDU (Lozhkov et al., 2024) for $10$B tokens, employing a batch size of $512$ and sequences of length $1024$ (see section G for additional hyperparameter details).
Results Table 2 reports the final loss achieved for all optimizers in our study, while figure 7 of the appendix plots the training curves. We observe that VeLO outperforms AdamW and $\mu$LOM, reaching the lowest final loss. This demonstrates that PyLO has the potential to deliver optimization performance improvements to PyTorch users.
5.3 Combining scheduler and decoupled weight decay
To showcase the interoperability of our library, we add learning rate scheduling and decoupled weight decay to our learned optimizer implementations in as little as $5$ lines of code. Although the primary objective of learned optimization is to eliminate the reliance on manually designed heuristics, we take the opportunity to investigate whether learning rate schedules and weight decay can still improve learned optimizer performance.
Effect of Learning Rate Scheduler We equip $\mu$LOM and VeLO with a cosine annealing LRS and tune the maximum learning rate (MaxLR). Figure 4 reports ViT and GPT training and accuracy curves for the best-performing MaxLR while training curves for more values are reported in the appendix (F). We observe that $\mu$LOM benefits substantially from explicit scheduling, extending stable training from 65k to 150k steps and reaching 71% Top-1 accuracy on ImageNet and improving loss in language model pre-training. VeLO shows minimal scheduling improvements, suggesting different internal adaptation mechanisms.
Effect of Weight decay We equip $\mu$LOM and VeLO with decoupled weight decay and tune the decay coefficient, $\lambda$. Figure 4 reports ViT and GPT training and accuracy curves for the best-performing $\lambda$ while training curves for more values are reported in the appendix (F). We observe that $\mu$LOM benefits from weight decay for training GPT but not VeLO. Neither optimizer benefits from weight decay for ViT tasks.
6 Conclusion
In conclusion, we have introduced PyLO, an open-source repository for state-of-the-art learned optimizers in PyTorch. PyLO breaks down the barriers to learned optimizers adoption by integrating directly with torch.optim.Optimizer and the HuggingFace ecosystem, allowing users to directly leverage learned optimizers in their existing code and easily download and share learned optimizer weights. For the best performance, PyLO implements a CUDA accelerated forward pass for the small_fc_lopt architecture, which we demonstrate has superior scaling properties and lower memory overhead than the original JAX implementation. When benchmarking the PyLO implementation of $\mu$LOM and VeLO against AdamW, we observed that VeLO can outperform AdamW on transformer language modeling and ViT tasks. We also ran ablations facilitated by PyLO’s PyTorch integration and found that learning rate schedules and weight decay can substantially improve the performance of certain learned optimizers.
Acknowledgments
We acknowledge support from FRQNT New Scholar [E.B.] and the FRQNT Doctoral (B2X) scholarship [B.T., A.M.]. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila.
Appendix A Comparison to Existing Learned Optimization Frameworks
Repository Overviews
Open-L2O (Chen et al., 2022) offers a comprehensive benchmarking suite designed primarily for evaluating learned optimizers (L2O) across a range of optimization problems. It includes both model-based and model-free L2O methods, facilitating reproducibility and fair comparisons but does not focus on practical deployment for real-world tasks.
Google’s learned_optimization (Metz et al., 2022a) is a research-oriented repository that provides a robust JAX-based framework for training, designing, evaluating, and applying learned optimizers. It extensively supports meta-training mechanisms, including Evolution Strategies (ES) and Predictive ES (PES), but lacks user-friendly interfaces for applying these optimizers in practical, production-oriented settings.
PyLO (ours) addresses critical barriers that have limited the adoption of learned optimizers in the broader machine learning community by providing an accessible, performance-focused PyTorch implementation. PyLO features seamless integration with PyTorch workflows, CUDA acceleration for computational efficiency, and HuggingFace ecosystem integration for easy sharing of optimizer weights and community-driven development.
Comparative Analysis
PyLO distinguishes itself from Open-L2O and Google’s learned_optimization in several key aspects:
  1. 1.
    Framework Compatibility: PyLO is built entirely in PyTorch, ensuring compatibility with the predominant framework used by practitioners, unlike Google’s learned_optimization, which is JAX-based.
  2. 2.
    Decoupled Meta-testing: PyLO exclusively provides optimized implementations for meta-testing without the complexity and overhead associated with meta-training code, unlike the other repositories.
  3. 3.
    HuggingFace Integration: PyLO integrates deeply with the HuggingFace Hub, facilitating standardized sharing, versioning, and benchmarking of task-specific optimizers, a capability not provided by the other repositories.
  4. 4.
    Computational Efficiency: By leveraging specialized CUDA kernels, PyLO significantly reduces optimizer step overhead, enabling practical deployment for large-scale models, a substantial improvement over naive implementations.
The strengths of PyLO position it uniquely as both researcher- and practitioner-friendly, enabling efficient exploration and practical use of state-of-the-art learned optimization techniques.
Table 3: Comparison of Learned Optimization Repositories
Repository HF PyTorch Decoupled Paper Github
Open-L2O (Chen et al., 2022) Open-L2O
Learned Optimization (Metz et al., 2022a) learned_optimization
PyLO (ours) In submission PyLO
Appendix B Extended challenges of L2O adoption for the Wider ML community
This section examines the critical barriers hindering the widespread adoption of learned optimizers within the broader machine-learning community. Despite their demonstrated potential for reduced training wall-clock time across various domains(Metz et al., 2022a; Chen et al., 2022), several structural and technical challenges have limited the integration of learned optimizers into standard machine learning workflows. The main barriers can be summarized as follows:
Several critical barriers have hindered the more widespread adoption of learned optimizers for deep learning tasks in our community: 1) overfocus on meta-learning in current frameworks, 2) absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, 4) the absence of learned optimizer benchmarking for popular machine learning tasks, and 5) the optimizer step overhead of learned optimizers.
  1. 1.
    The overfocus on meta-learning in current frameworks.
  2. 2.
    The absence of a PyTorch implementation for the most performant learned optimizers available.
  3. 3.
    The lack of a standardized way to effectively share learned optimizer weights.
  4. 4.
    The absence of learned optimizer benchmarking for popular machine learning tasks.
  5. 5.
    The optimizer step overhead of learned optimizers.
Overfocus on meta-learning
A primary obstacle to L2O adoption stems from the disparity between current learned optimization frameworks and established training paradigms. Existing libraries, such as the JAX (Bradbury et al., 2018)-based learned optimization framework (Metz et al., 2022a, b), prioritize meta-training capabilities over practical deployment. While these frameworks do provide mechanisms for utilizing pre-trained learned optimizers, the interfaces diverge significantly from those familiar to practitioners (Wightman, 2019; Wolf et al., 2020; von Platen et al., 2022). The evaluation and application of learned optimizers appear as secondary objectives, with insufficient emphasis on seamless integration with conventional training pipelines. This architectural choice highlights the necessity of decoupling meta-learning and meta-testing phases, particularly for researchers and practitioners who seek to leverage pre-trained optimization algorithms without engaging in the complexities of meta-training.
Pytorch v.s. JAX
A second significant barrier relates to the underlying deep learning frameworks. While JAX (Bradbury et al., 2018) offers powerful functional programming capabilities and high flexibility for research workflows, it lacks the mature ecosystem and widespread adoption of more established frameworks. In contrast, PyTorch (Paszke et al., 2019) has developed into the dominant deep learning framework with extensive community support (He, 2019) and a rich ecosystem of pre-trained models and resources (Wightman, 2019; von Platen et al., 2022; Wolf et al., 2020). This prevalence stems partly from PyTorch’s object-oriented architecture and modular design principles, which facilitate extension, ease-of-use, and customization. The absence of a comprehensive learned optimizer library that conforms to PyTorch’s optimizer interface presents a substantial obstacle, as it necessitates significant codebase modifications for implementation and experimentation. Consequently, there exists a pressing need for a modular L2O framework that seamlessly integrates with PyTorch’s optimizer class hierarchy. Such integration, particularly when combined with standardized weight serialization and distribution through established repositories, would substantially accelerate both the circulation and scaling of learned optimization techniques.
Absence of standardized benchmarking
A third critical limitation concerns evaluation methodologies and benchmarking procedures. Current L2O framework Metz et al. (2022a) , typically evaluate optimization algorithms across an extensive but potentially misaligned task distribution. These evaluations predominantly focus on training loss trajectories rather than generalization performance metrics that are of primary interest to practitioners. Moreover, recent works like VeLO(Metz et al., 2022b) omit crucial experimental details, including baseline hyperparameter configurations , impeding reproducibility and comparative analysis(Rezk et al., 2023). This deficiency underscores the necessity for standardized evaluation frameworks that enable systematic comparison of conventional and learned optimizers on practically relevant tasks. Such frameworks should prioritize contemporary challenges in vision and language model pre-training rather than theoretical convex optimization problems that, while analytically tractable, often fail to capture the complexities of modern neural network training dynamics.
Sharing learned optimizer weights
Pre-trained models for language understanding (DeepSeek-AI et al., 2025; Touvron et al., 2023), image classification (Dosovitskiy et al., 2020; Radford et al., 2021), semantic segmentation (Kirillov et al., 2023), and a number of other tasks are routinely shared via platforms such as HuggingFace Hub (HuggingFace, ), enabling practitioners to leverage computational investments made by well-resourced institutions. This collaborative open-model ecosystem has accelerated innovation and reduced redundancy in these domains through its expansion of access to state-of-the-art models. Despite this revolution in model sharing, we lack seamless sharing in learned optimizer weights. We suggest that foundational Learned Optimizer (LO) models can and should be integrated into the community following similar distribution mechanisms (e.g., HuggingFace) to those established for neural network weights. Our objective is to facilitate seamless integration of learned optimizers into existing PyTorch (Paszke et al., 2019) workflows with minimal friction, addressing the significant gap in accessibility for the predominant deep learning framework used by practitioners. Furthermore, by establishing infrastructure for distributing optimizer weights, we aim to catalyze a parallel ecosystem where learned optimizer rules can be shared and deployed across diverse problem domains.
Per-step overhead of learned optimizers
A significant obstacle to the widespread adoption of learned optimizers is the computational overhead they potentially introduce to already resource-intensive training setups. Deep learning workflows operate under strict time and computational constraints, with practitioners demonstrating marked reluctance to try new optimization strategies that might increase training durations, regardless of potential convergence benefits. This challenge necessitates meticulous implementation of learned optimization algorithms to ensure competitive computational efficiency. Our approach addresses this critical constraint through custom CUDA implementations that substantially reduce optimizer step time compared to native PyTorch implementations. By implementing the learned optimizer update as a neural network pass directly in CUDA and strategically storing intermediate activations in registers, we minimize the need to write activations to global GPU memory. This optimization achieves substantial reductions in wall-clock time for optimization steps and render learned optimizers practically viable within existing computational budgets, removing a significant barrier to their adoption by the broader machine-learning community.
Appendix C Extended Code design
We implement a modular architecture partitioned into three main components to realize our project vision. This design philosophy prioritizes both usability for practitioners and extensibility for researchers developing novel learned optimizers.
Optimization Module (pylo.optim) This module facilitates critical state management functions necessary for learned optimization. These include maintaining parameter-specific accumulators, computing parameter features for optimizer forward passes, executing update steps with configurable learning rate scaling, and supporting standard PyTorch optimizer functionality such as state dictionaries for resuming. This implementation allows practitioners to leverage advanced learned optimization techniques without significant modifications to existing training setups.
Meta-Model Architectures (pylo.models) encapsulates the parameters of the learned optimizer and its forward pass. It is also fully integrated with HuggingFaceHub (HuggingFace, ), allowing users to download existing optimizers from the hub and providing a mechanism for weight distribution and versioning of future optimizers. Through HF integration, we facilitate community-driven improvement cycles and allow researchers to easily leverage our CUDA-accelerated implementation for their own benchmarking.
CUDA Acceleration (pylo.csrc) The computational demands of learned optimizers present significant implementation challenges. Unlike traditional optimization algorithms that apply simple, closed-form update rules, learned optimizers require evaluating a small multilayer perceptron (MLP) for each parameter in the optimizee. This can introduce substantial computational overhead for large models, creating a critical bottleneck that limits practical deployment. Our analysis revealed that standard PyTorch neural network modules are inadequately optimized for this specific use case. These modules are primarily designed for batched operations on image or language data rather than the unique computational pattern of learned optimizers, where many small MLPs must be evaluated in parallel across millions of parameters.
Drawing inspiration from the architecture-aware optimizations (Müller et al., 2021; Dao et al., 2022) we implemented a specialized CUDA kernel for the small_fc_lopt learned optimizer. This implementation strategically leverages the GPU memory hierarchy to address the primary bottleneck: memory bandwidth rather than computational throughput. Our CUDA implementation employs two key optimization strategies: (1) Register-Based Computation and (2) On-the-Fly Feature Computation. The first utilizes GPU registers to store intermediate activations during forward propagation through the optimizer’s MLP. This approach minimizes high-latency global memory accesses by keeping frequently accessed data in the fastest available memory tier. The second avoids storing normalized features in global memory, recalculating them when needed, and trading redundant computation for reduced memory bandwidth.
Appendix D Extended Experimental Results
D.1 Benchmarking LO step times in PyLO
Model Optimizer Batch Size Samples per sec Step time (ms) Forward time (ms) Backward time (ms) Optimizer step time(ms) Num parameters(M) Vit-B-16 AdamW 32.00 524.09 60.55 17.52 38.13 4.90 86.57 Adafactor 425.51 74.69 18.99 small_fc_lopt (naive) 39.36 812.51 756.80 small_fc_lopt (CUDA) 205.59 155.16 99.59 Vit-S-16 AdamW 32.00 1,116.12 28.17 7.65 18.27 2.25 22.05 Adafactor 711.97 44.47 18.53 small_fc_lopt (naive) 109.03 293.01 266.90 small_fc_lopt (CUDA) 382.73 83.28 57.45 Vit-L-16 AdamW 32.00 175.95 181.36 52.18 113.89 15.28 304.33 Adafactor 156.64 203.34 37.23 small_fc_lopt (naive) 11.64 2,748.52 2,582.30 small_fc_lopt (CUDA) 70.69 451.69 285.80 GPT2-125 m AdamW 4.00 17.80 224.72 73.20 144.24 7.29 125.26 Adafactor 17.02 235.07 17.76 small_fc_lopt (naive) 3.38 1,182.33 964.89 small_fc_lopt (CUDA) 11.71 341.46 124.09 GPT2-355 m AdamW 4.00 6.56 609.83 197.41 392.29 20.12 355.92 Adafactor 6.40 624.64 35.11 small_fc_lopt (naive) 1.16 3,461.96 2,872.17 small_fc_lopt (CUDA) 4.40 908.69 319.14 GPT2-1B AdamW 4.00 2.88 1,387.51 442.49 896.81 48.21 912.95 Adafactor 2.87 1,395.03 54.08 small_fc_lopt (naive) OOM OOM OOM small_fc_lopt (CUDA) 1.98 2,025.26 681.78
Table 4: Breakdown of time spent per training step, averaged over 40 steps. The CUDA-based implementation reduces overhead from the learned optimizer by minimizing memory usage and optimizer time, ultimately achieving competitive throughput compared to traditional optimizers.
We extend our analysis beyond the baseline results to examine per-step computational overhead across varying model scales representative of contemporary pretraining workloads. Table 4 presents a comprehensive timing breakdown across forward pass, backward pass, and optimizer step durations for both vision and language model architectures. Traditional optimizers including AdamW (Loshchilov & Hutter, 2019) and Adafactor (Shazeer & Stern, 2018) exhibit minimal computational overhead relative to the forward and backward passes. Our CUDA kernel implementations demonstrate substantial reductions in learned optimizer overhead, with improvements that scale favorably with model size. Notably, for ViT-L-16, our small_fc_lopt(CUDA) implementation achieves a 9× reduction in optimizer step time compared to the naive implementation. This performance advantage becomes increasingly pronounced at scale. Consistent with our earlier findings regarding memory limitations—where JAX implementations encounter out-of-memory errors for 1B parameter language models—the naive learned optimizer implementation similarly fails at this scale. Our CUDA implementation addresses these scalability constraints while maintaining the computational efficiency gains observed with increasing batch sizes, making learned optimization viable for large-scale pretraining scenarios.
(a) Width
(b) Depth
Figure 5: Step time for deeper and wider MLPs. Increasing the width and depth of MLPs provides a systematic way to measure the impact of larger and more tensors on optimizer step time. Subfigure (a) reports optimizer step time as the hidden size of the 5-layer optimizee MLP is increased, while Subfigure (b) reports step time as the depth of a 256 hidden-layer MLP is increased. We observe that in each case, the CUDA implementation reduces the optimizer step time by more than an order of magnitude.
Scaling analysis with MLP: To systematically characterize the scaling behavior of our implementation, we conduct controlled experiments using synthetic MLP architectures with varying architectural parameters. Figure 5 presents two complementary scaling analyses that isolate the effects of network width and depth on optimizer computational overhead. Figure 5(a) examines the relationship between hidden layer dimensionality and optimization time for a fixed 5-layer MLP architecture. The results reveal a stark performance disparity: the naive implementation of small_fc_lopt exhibits optimization times an order of magnitude greater than our CUDA implementation and encounters memory limitations beyond hidden dimensions of 256. While our CUDA implementation remains slower than traditional optimizers (AdamW and Adafactor), it demonstrates favorable scaling characteristics that maintain computational tractability across the tested range.
Complementary analysis in Figure 5(b) evaluates depth scaling behavior using MLPs with fixed width of 256 units but varying layer counts (depth). The naive implementation consistently operates in a computationally prohibitive regime, requiring over 400ms for optimization steps, while our CUDA implementation maintains optimization times below 50ms—approaching the performance envelope of traditional optimizers despite the inherent complexity of learned optimization. These scaling experiments demonstrate that our CUDA implementation successfully addresses the computational bottlenecks that render naive learned optimizer implementation impractical for realistic model architectures.
(a) ViT small patch 16
(b) ViT base patch 16
(c) ViT large patch 16
Figure 6: Vision Transformer training dynamics. Training loss curves for Vision Transformers trained on ImageNet-1k classification across three model configurations: (a) ViT-small with patch size 16, (b) ViT-base with patch size 16, and (c) ViT-large with patch size 16. We compare AdamW with cosine learning rate scheduling (blue), VeLO learned optimizer (orange), and $\mu$LO learned optimizer (black) over 150,000 training steps with batch size 4,096. $\mu$LO exhibits early training instability across all scales, while VeLO demonstrates competitive or superior convergence properties, particularly evident in the base model configuration where it achieves lower final loss than the AdamW baseline.
Appendix E Training Curves for the Experiments
This section presents a comprehensive empirical evaluation comparing learned optimizers VeLO (Metz et al., 2022b) and $\mu$LO (Thérien et al., 2024) against the standard AdamW optimizer (Loshchilov & Hutter, 2019; Kingma & Ba, 2017) with cosine learning rate scheduling (Loshchilov & Hutter, 2017). Our evaluation spans Vision Transformer architectures of varying scales and GPT-2 models, extending beyond the main paper’s validation accuracy results for ViT-Base/16 and final loss metrics for GPT-2 to provide detailed training dynamics across multiple configurations.
Vision Transformer Experiments: We evaluate three ViT configurations—Small, Base, and Large with patch size 16—on ImageNet-1k classification tasks. All configurations maintain consistent training protocols with 150,000 optimization steps and batch sizes of 4,096 across model variants, using hyperparameters detailed in Table 6. The extended evaluation presented in Figure 6 reveals critical scale-dependent optimization characteristics. MuLO demonstrates pronounced early divergence patterns across both small and large model configurations, reinforcing our hypothesis that its constrained meta-training horizon limits generalization to extended optimization trajectories. In contrast, VeLO exhibits remarkable consistency in convergence properties across the full spectrum of model scales examined, maintaining stable training dynamics while achieving competitive performance metrics relative to the AdamW baseline.
Language Modeling Experiments For GPT-2 pretraining on FineWeb-EDU data, we train models for 19,073 steps with batch size 512 and sequence length 1024. Complete hyperparameters are presented in Table 8, with full training curves shown in Figure 7. VeLO demonstrates competitive performance against AdamW with cosine scheduling, while MuLO exhibits significantly degraded performance across the extended training horizon, consistent with the stability patterns observed in vision tasks.
(a) 355M LM
Figure 7: GPT-2 pretraining performance. We pre-train decoder-only transformers of different sizes on a causal language modeling objective. We estimate gradients from 512 sequences of length 1024, resulting in a $\sim$0.5M token batch size per step. We train all models for 10B tokens of FineWeb-EDU data.
Appendix F Extended Ablation Studies
This section presents comprehensive training curves across all hyperparameter configurations examined in our weight decay and cosine annealing ablations. Our analysis encompasses systematic sweeps of weight decay coefficients and maximum learning rates for cosine annealing schedules.
Vision Transformer Ablations Figures 8 and 9 demonstrate the variation of training dynamics with varying maximum learning rates and weight decay values for the ViT-B/16 model. The results reveal that VeLO exhibits minimal sensitivity to weight decay and scheduler modifications in this experimental setup, suggesting robust inherent regularization properties. Conversely, MuLO shows improved training horizon stability when augmented with cosine scheduling, enabling successful completion of the full 150,000 training steps. Weight decay modifications yield marginal improvements for MuLO across the evaluated parameter ranges.
Language Modeling Ablations
The language modeling experiments, presented in Figures 10 and 11, demonstrate more pronounced sensitivity to weight decay and scheduling compared to vision tasks. VeLO shows measurable performance improvements when augmented with weight decay and scheduling for specific parameter configurations. Similarly, MuLO benefits substantially from both weight decay regularization and cosine scheduling in the language modeling domain, suggesting that learned optimizers may benefit from scheduling and weight decay.
Architectural Configuration Effects
Figure 12 presents an additional ablation examining the impact of separate versus concatenated query-key-value (QKV) matrices on the optimizee’s training loss. We observe that separating the QKV matrices in attention blocks improves performance across VeLO and $\mu$LOM.
(a) VeLO Decoupled Weight decay ablation
(b) VeLO Cosine annealing Ablation
Figure 8: Training ViT-B-16 with VeLO on Imagenet 1k with decoupled weight decay and cosine annealing. We see no improvement in using weight decay or Scheduler for VeLO in this setup
(a) $\mu$LO Decoupled Weight decay ablation
(b) $\mu$LO Cosine annealing Ablation
Figure 9: Training ViT-B/16 with $\mu$LO on Imagenet1k with decoupled weight decay and cosine annealing. We see that the training horizon of $\mu$LO is improved with using a scheduler that helps it to run up to 150000 training steps. We also see that weight decay did not have substantial improvement
(a) VeLO Decoupled Weight decay ablation
(b) VeLO Cosine annealing Ablation
Figure 10: GPT2 Pre-training VeLO ablations: decoupled weight decay and cosine annealing. We pre-train decoder-only transformers of different sizes on a causal language modeling objective. We estimate gradients from $512$ sequences of length $1024$, resulting in a $\sim 0.5$ M token batch size per step. We train all models for $10B$ tokens of FineWeb-EDU data.
(a) $\mu$LO Decoupled Weight decay ablation
(b) $\mu$LO Cosine annealing Ablation
(c) $\mu$LO LR Tuning Ablation
Figure 11: GPT2 Pre-training VeLO ablations: decoupled weight decay, LR-tuning, and cosine annealing. We pre-train decoder-only transformers of different sizes on a causal language modeling objective. We estimate gradients from $512$ sequences of length $1024$, resulting in a $\sim 0.5$ M token batch size per step. We train all models for $10B$ tokens of FineWeb-EDU data.
(a) VeLO
(b) $\mu$LO
Figure 12: Separate QKV weight ablation. We ablate using concatenated or separate QKV matrices in attention layers. From the perspective of the learned optimizer, these two settings will be treated differently. We observe that performance improves for both $\mu$LOM and VeLO when QKV matrices are separated.
Appendix G Hyperparameters for Experiments
Table 5: Vision Transformer Training Hyperparameters.
Description Value
Model Architecture
Model ViT-Base/16
Image Size $224\times 224$
Patch Size $16\times 16$
Training Configuration
Batch Size $4096$
Training Epochs $480$
Optimizer (for baseline) AdamW
Learning Rate ($\eta$) $4\times 10^{-3}$
LR Scheduler Cosine
Warmup Epochs $32$
Weight Decay $0.03$
Gradient Clipping $1.0$
Data Augmentation
Crop Percentage $0.95$
Random Horizontal Flip $0.5$
Mixup $0.1$
CutMix $1.0$
AutoAugment rand-m7-mstd0.5
Regularization
Dropout Rate $0.1$
Stochastic Depth $0.1$
Table 6: Vision Transformer Model Variants Architecture.
Description Value
ViT-Small/16
Parameters $22.05$M
Hidden Dimension ($d_{\text{model}}$) $384$
MLP Hidden Dimension $1536$
Number of Heads $6$
Number of Layers $12$
Patch Embedding Dim $384$
ViT-Base/16
Parameters $86.57$M
Hidden Dimension ($d_{\text{model}}$) $768$
MLP Hidden Dimension $3072$
Number of Heads $12$
Number of Layers $12$
Patch Embedding Dim $768$
ViT-Large/16
Parameters $307.40$M
Hidden Dimension ($d_{\text{model}}$) $1024$
MLP Hidden Dimension $4096$
Number of Heads $16$
Number of Layers $24$
Patch Embedding Dim $1024$
Table 7: GPT-2 Training Hyperparameters.
Description Value
Model Configuration
Base Model GPT-2
Architecture Type Decoder-only Transformer
Attention Mechanism separate_kqv
Sequence Length ($T$) $1024$
Vocabulary Size $50257$
Training Configuration
Batch Size $512$
Training Iterations $19073$
Optimizer (for baseline) AdamW
Learning Rate Scheduler (for baseline) Cosine
Warmup Iterations $381$
Regularization
Attention Dropout $0.1$
Embedding Dropout $0.1$
Residual Dropout $0.1$
Initialization
Muon Initialization Enabled
Weight Initialization Standard GPT-2
Table 8: GPT-2 Model Variants Architecture.
Description Value
GPT-2 Small
Parameters $\sim 36$M
Hidden Dimension ($d_{\text{model}}$) $768$
Number of Heads $12$
Number of Layers $12$
Head Dimension $64$
FFN Hidden Dimension $3072$
GPT-2 Medium
Parameters $\sim 345$M
Hidden Dimension ($d_{\text{model}}$) $1024$
Number of Heads $16$
Number of Layers $24$
Head Dimension $64$
FFN Hidden Dimension $4096$
GPT-2 Large
Parameters $\sim 762$M
Hidden Dimension ($d_{\text{model}}$) $2048$
Number of Heads $32$
Number of Layers $16$
Head Dimension $64$
FFN Hidden Dimension $8192$

### v2

@@
-PyLO: Towards Accessible Learned Optimizers in PyTorch
+
 Abstract
 Learned optimizers have been an active research topic over the past decade, with increasing progress toward practical, general-purpose optimizers that can serve as drop-in replacements for widely used methods like Adam.
-However, recent advances—such as VeLO, which was meta-trained for 4000 TPU-months, remain largely inaccessible to the broader community, in part due to their reliance on JAX and the absence of user-friendly packages for applying the optimizers after meta-training.
-To address this gap, we introduce PyLO, a PyTorch-based library that brings learned optimizers to the broader machine learning community through familiar, widely adopted workflows.
-Unlike prior work focused on synthetic or convex tasks, our emphasis is on applying learned optimization to real-world large-scale pre-training tasks.
-Our release includes a CUDA-accelerated version of the small_fc_lopt learned optimizer architecture from (Metz et al., 2022a), delivering substantial speedups—from $39.36$ to $205.59$ samples/sec throughput for training ViT-B/16 with batch size $32$.
-PyLO also allows us to easily combine learned optimizers with existing optimization tools such as learning rate schedules and weight decay.
-When doing so, we find that learned optimizers can substantially benefit.
-Our code is available at https://github.com/Belilovsky-Lab/pylo
-Machine Learning, ICML
+However, recent advances such as VeLO, which was meta-trained for 4000 TPU-months, remain largely inaccessible to the broader community, in part due to their reliance on JAX and the absence of user-friendly packages for independently using the optimizers after meta-training.
+To address this gap, we introduce PyLO, a PyTorch-based library that brings learned optimizers to the remaining $\approx 80\%$ of machine learning community via the familiar torch.optim.Optimizer interface.
+Unlike prior work focused on limited-scale academic tasks, our emphasis is on applying learned optimization to real-world large-scale pre-training tasks.
+Our systems contribution includes CUDA-accelerated implementations of the small_fc_loptMetz et al.
+(2022a) and VeLOMetz et al.
+(2022b) learned optimizers, achieving substantial performance gains, with training throughput on ViT-B/16 (batch size 32) increasing from 39.36 and 49.73 to 205.59 and 191.18 samples per second, respectively.
+PyLO has the versatility that allows us to easily combine learned optimizers with existing optimization tools, such as learning rate schedules and weight decay.
+When doing so, we discover that learned optimizers can substantially benefit from it.
+Our code is available at https://github.com/Belilovsky-Lab/pylo .
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
 Figure 1: Training step timing breakdown for ViT-B/16 on a single A100 GPU.
-We measure the forward (blue), backward (orange), and optimizer (other colors) times per training step.
-We observe that the CUDA-accelerated learned optimizer step shows substantial improvements over the naive implementation (red) and is competitive with adafactor (purple).
+We measure the forward, backward, and optimizer times per training step.
+We observe that the CUDA-accelerated learned optimizer steps (cyan, green) show substantial improvements over the naive implementations (yellow, red).
 In all cases, as the batch size is increased, the relative overhead of the optimizer shrinks.
@@
 By addressing key usability challenges, PyLO provides seamless access to meta-learned optimization techniques through three core features: (1) automatic weight loading from Hugging Face Hub, (2) a familiar PyTorch-style optimizer interface, and (3) accelerated CUDA kernel support.
-The library bridges the gap between advanced meta-learning research and practical machine-learning applications, enabling researchers and practitioners to easily leverage state-of-the-art learned optimization techniques in Pytorch.
-Learned optimization (LO) (Hochreiter et al., 2001; Andrychowicz et al., 2016) represents a promising yet underutilized direction for advancing optimization algorithms in machine learning.
+The library bridges the gap between advanced meta-learning research (google/learned_optimizationMetz et al.
+(2022a)) and practical machine-learning applications, enabling researchers and practitioners to easily leverage state-of-the-art learned optimization techniques in PyTorch.
+Learned optimization (LO) (Hochreiter et al., 2001; Andrychowicz et al., 2016) represents a promising yet underexplored direction for advancing optimization algorithms in machine learning.
 Training contemporary neural networks requires solving highly non-convex optimization problems for which theory neither guarantees convergence to a globally optimal solution, nor convergence at the optimal rate.
-Therefore, despite the purportedly strong performance of hand-designed optimizers such as Adam (Kingma & Ba, 2017), for many tasks of interest, there may be substantial room for improvement.
-Therefore, data-driven optimization algorithms have the potential to make substantial improvements.
-However, even the most performant publically available learned optimizer to date, VeLO (Metz et al., 2022b), has seldom been adopted by deep learning practitioners.
-Several critical barriers have hindered the more widespread adoption of learned optimizers: 1) an overfocus on meta-learning in current learned optimization frameworks, 2) the absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, 4) the absence of learned optimizer benchmarking for popular machine learning tasks, and 5) learned optimizer step overhead.
+Therefore, despite the purportedly strong performance of hand-designed optimizers such as Adam (Kingma & Ba, 2017), for many tasks of interest, there may be substantial room for improvement through learning to optimize.
+Indeed, VeLO Metz et al.
+(2022b), the most performant publicly available learned optimizer to date, was shown to outperform a well-tuned NadamW and schedule baseline optimizer without hyperparameter tuning.
+Despite its strong performance, however, VeLO is seldom used by our community today.
+Several critical barriers have hindered the more widespread adoption of learned optimizers for deep learning applications: 1) an overfocus on meta-learning in current learned optimization frameworks, 2) the absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, and 4) learned optimizer step overhead.
 In what follows, we take a step towards improving the adoption of learned optimizers for deep learning by providing a PyTorch implementation of state-of-the-art learned optimizers that emphasizes their performance and accessibility for practical machine learning problems.
@@
   1. 1.
-    We provide a modular open-source implementation of state-of-the-art learned optimizers in PyTorch which seamlessly integrates with torch.optim.Optimizer and the Huggingface ecosystem to enable easy usage within existing code and standardized sharing of task-specific learned optimizer weights.
+    We provide a modular open-source implementation of state-of-the-art learned optimizers in PyTorch, which seamlessly integrates with torch.optim.Optimizer and the Huggingface ecosystem to easily integrate with existing code and facilitate standardized sharing of task-specific learned optimizer weights (Figure 2).
   2. 2.
-    We provide a CUDA-accelerated implementation of small_fc_lopt’s optimizer step which substantially reduces memory overhead and increases occupancy (GPU utilization as in Figure 1).
+    We provide a CUDA-accelerated implementation of small_fc_lopt and VeLO’s optimizer step, which substantially reduces memory overhead and increases occupancy (GPU utilization as in Figure 1).
   3. 3.
-    We benchmark optimizer step times and performance in PyLO for popular image classification and language modeling workloads, showing that our implementation scales to large workloads and improves on the existing JAX implementation.
+[…]
   4. 4.
-    Our flexible framework makes it easy to study the effect of weight decay and learning rate schedules on learned optimizer performance and, when doing so, we find that making these simple additions can substantially improve performance in some cases.
-2 Project vision
+    We demonstrate that learned optimizer step times can further be reduced in a data-parallel training setting by distributing the optimizer step across devices.
+  5. 5.
+    Our flexible framework makes it easy to study the effect of weight decay and learning rate schedules on learned optimizer performance by leveraging the existing Pytorch API and, when doing so, we find that making these simple additions can substantially improve performance in some cases.
+2 Challenges of L2O adoption for the Wider deep learning community
+This section reviews existing open-source learned optimization repositories and reflects on the critical barriers hindering the widespread adoption of learned optimizers within the broader deep learning community, highlighting pitfalls in existing work.
+Table 1: Comparison of Open-source Learned Optimization Repositories.
+We list the current open-source repositories for learned optimization and list their status relative to the difficulties of L2O adoption from section $2.$ A Decoupled framework separates learned optimizer implementation from meta-training and meta-testing code.
+Repository Decoupled PyTorch HuggingFace CUDA Kernel Paper Github Open-L2O Chen et al. (2022) Open-L2O Learned Optimization Metz et al. (2022a) Learned Optimization PyLO (ours) In submission PyLO
+2.1 Existing Open source libraries
+Open-L2O Chen et al.
+(2022) offers a comprehensive benchmarking suite designed primarily for evaluating learned optimizers (L2O) across a range of different optimization problems: convex functions, non-convex functions, minimax functions, and some neural network training.
+It includes both model-based and model-free L2O methods, facilitating reproducibility and fair comparisons, but only includes a limited evaluation of the learned optimizer for practical deep learning tasks.
+In contrast, PyLO’s main focus is on learned optimization for deep learning.
+Google’s learned_optimization Metz et al.
+(2022a) is a research-oriented repository written in JAX for training, designing, evaluating, and applying learned optimizers.
+It supports an extensive set of features for meta-training optimizers: many gradient estimation algorithms, abstractions for different types of tasks, various task modifiers, etc.
+While the repository also provides functionality for applying pre-trained optimizers to new tasks via the Optax interface DeepMind et al.
+(2020), its main focus is clearly to provide meta-training functionality.
+2.2 Challenges to adoptions
+Overfocus on meta-learning: Existing libraries, such as the Learned Optimization (Metz et al., 2022a; b) and Open L2O Chen et al.
+(2022), prioritize meta-training functionality over precisely focusing on the practical deployment of learned optimizers.
+This leads to a large proportion of the code being unnecessary for applying learned optimizers to new tasks.
+This can represent a barrier to entry, particularly for researchers and practitioners who seek to leverage pre-trained optimization algorithms without engaging in the complexities of meta-training (similar to using a pretrained transformer from huggingface/transformersWolf et al.
+(2020)).
+This highlights the necessity of decoupling learned optimizer implementations from their meta-training and meta-testing code.
+PyTorch v.s.
+JAX: A third barrier relates to the underlying deep learning framework.
+While JAX (Bradbury et al., 2018) offers powerful functional programming capabilities and strong performance from compilation, it lacks the widespread community adoption of PyTorch (Paszke et al., 2019).
+Beyond having more users, widespread community adoption also comes with a more mature and feature-full open source ecosystem He (2019); Wightman (2019); von Platen et al.
+(2022); Wolf et al.
+(2020).
+With so much demand for compatibility with PyTorch, there exists a pressing need for a modular L2O framework that seamlessly integrates with PyTorch’s optimizer interface.
+Currently, Google’s Learned Optimization does not provide any compatibility with PyTorch.
+While Open-L2O does support PyTorch, it lacks many other crucial aspects for adoption (see Table 1).
+Sharing learned optimizer weights: Being able to effectively share optimizer weights and document their training distribution and evaluation performance is essential for building an effective open source ecosystem around learned optimizers.
+However, no such mechanism currently exists.
+In contrast, pre-trained models for language understanding (DeepSeek-AI et al., 2025; Touvron et al., 2023), image classification (Dosovitskiy et al., 2020; Radford et al., 2021), semantic segmentation (Kirillov et al., 2023), and a number of other tasks are routinely shared via platforms such as HuggingFace Hub (HuggingFace, ), enabling practitioners to leverage computational investments made by well-resourced institutions.
+This collaborative open-model ecosystem has accelerated innovation and reduced redundancy in these domains through its expansion of access to state-of-the-art models.
+Despite this revolution in model sharing, we lack seamless sharing in learned optimizer weights.
+We believe that Learned Optimizer (LO) models can and should be integrated into the community through the HuggingFace ecosystem.
+Current open-source L2O frameworks, Learned Optimization and OpenL2O, do not provide any mechanism for sharing learned optimizer weights.
+Per-step overhead of learned optimizers: A significant obstacle to the widespread adoption of learned optimizers is the computational overhead they can introduce to already resource-intensive training setups.
+Deep learning workflows operate under strict time and computational constraints, with practitioners demonstrating marked reluctance to trying new optimization strategies that might increase training durations, regardless of potential convergence benefits.
+This challenge necessitates meticulous implementation of learned optimization algorithms to ensure competitive computational efficiency, which can be achieved by writing custom CUDA kernels.
+3 Project goal
 With our goal of improving the adoption of learned optimizers within the deep learning community and enhancing the open-source ecosystem around them, we design PyLO with the following principles in mind:
@@
 As a result, PyLO is simpler to use and easier to understand.
-Interoperability Through strict adherence to the PyTorch optimizer (torch.optim.Optimizer) interface specification, PyLO ensures seamless integration with existing training pipelines, enabling practitioners to adopt learned optimizers with minimal modification to existing code.
-It also helps to use external learning rate schedulers and decoupled weight decay.
-3 Code design
-We implement a modular architecture partitioned into three main components to realize our project vision.
-This design philosophy prioritizes both usability for practitioners and extensibility for researchers developing novel learned optimizers.
-Optimization Module (pylo.optim) This module facilitates critical state management functions necessary for learned optimization.
-These include maintaining parameter-specific accumulators, computing parameter features for optimizer forward passes, executing update steps with configurable learning rate scaling, and supporting standard PyTorch optimizer functionality such as state dictionaries for resuming.
-This implementation allows practitioners to leverage advanced learned optimization techniques without significant modifications to existing training setups.
-Meta-Model Architectures (pylo.models) encapsulates the parameters of the learned optimizer and its forward pass.
-It is also fully integrated with HuggingFaceHub (HuggingFace, ), allowing users to download existing optimizers from the hub and providing a mechanism for weight distribution and versioning of future optimizers.
-Through HF integration, we facilitate community-driven improvement cycles and allow researchers to easily leverage our CUDA-accelerated implementation for their own benchmarking.
-CUDA Acceleration (pylo.csrc) The unique computational demands of learned optimizers present challenges that default PyTorch is unequipped to handle.
-Therefore, we implemented a specialized CUDA kernel for the small_fc_lopt learned optimizer.
-This implementation strategically leverages the GPU memory hierarchy to address the primary bottleneck: memory bandwidth rather than computational throughput.
-Our CUDA implementation employs two key optimization strategies: (1) Register-Based Computation and (2) On-the-Fly Feature Computation.
-The first utilizes GPU registers to store intermediate activations during forward propagation through the optimizer’s MLP.
-This approach minimizes high-latency global memory accesses by keeping frequently accessed data in the fastest available memory tier.
-The second avoids storing normalized features in global memory, recalculating them when needed, and trading redundant computation for reduced memory bandwidth.
-Figure 3: Learned Optimizer step time scaling plot.
-We compare the optimizer step time for CUDA and JAX implementations of small_fc_lopt.
-We observe that our CUDA implementation allows scaling to larger models and substantially reduces step time.
-4 Benchmarking LO step times in PyLO
-In this section, we establish the per-step overhead of learned optimizers for common pre-training workloads and showcase the improved scaling behavior of PyLO’s CUDA-accelerated learned optimizer steps.
-Our goal is to report learned optimizer overhead relative to hand-designed optimizers for practical tasks and to illustrate how the overhead scales as the model size is increased.
-To this end, we benchmark and record the forward, backward, and optimizer step times for training ViT-B/16 and a $410$M parameter language model on a single $80$GB A100 GPU.
-Learned Optimizer overhead shrinks as batch size increases Figure 1 reports timings at different batch sizes for ViT-B/16.
-We observe that as batch size is increased the overhead of the learned optimizer’s step compared to the forward and backward pass shrinks.
-This suggests that large-scale training with long per-step times may be a suitable workload for learned optimizers, where the cost of applying the optimizer can be amortized.
-Table 1 compares timing results across different optimizers for ViT and GPT.
-We observe that the per-step time small_fc_lopt (CUDA) increases due to the large parameter count of the language model but that using our CUDA kernels still leads to a remarkable improvement over the naive implementation.
-We note that it also halves the optimizer step time of the original JAX implementation.
-Scaling behavior of our CUDA implementation Figure 3 reports the scaling behavior w.r.t.
-transformer size of different small_fc_lopt (we use hidden size $32$ as in (Thérien et al., 2024)) implementations: CUDA and JAX.
-JAX corresponds to the original implementation of (Metz et al., 2022a) which uses the JAX compiler, while CUDA is the fastest implementation that we make available in PyLO.
-We observe that our CUDA implementation results in a substantial memory savings as it can optimize the large model size while the JAX implementation fail for the 1B parameter transformer.
-Moreover, our CUDA implementation also outperforms the JAX compiler by 2x, which corresponds to a substantial reduction in optimizer overhead.
-Model Optimizer BS/SL Fwd (ms) Bwd (ms) Opt step (ms) #Params (M)
-ViT-B/16 AdamW 32/197 17.52 38.13 4.90 86.57
-Adafactor 18.99
-small_fc_lopt (naive) 756.80
-small_fc_lopt (CUDA) 99.59
-GPT2-355M AdamW 4/1024 197.41 392.29 20.12 355.92
-Adafactor 35.11
-small_fc_lopt (naive) 2872.17
-small_fc_lopt (CUDA) 319.14
-Table 1: Timing results across different optimizers for ViT and GPT. We report timings for realistic batch size (BS) and sequence length (SL) which fit on a single A100 GPU.
-5 Demonstrating the real world use of PyLO
-In the following section, we evaluate and compare our PyTorch implementation of VeLO (Metz et al., 2022b) and $\mu$LOM (Thérien et al., 2024) to the performance of tuned AdamW with a cosine annealing schedule.
-Our goal is to establish the performance of readily available learned optimizers in PyLO for these practical vision and language pre-training tasks.
-(a) LM $\mu$LO
-(b) LM VeLO
-(c) ViT $\mu$LO
-(d) ViT VeLO
-Figure 4: Weight decay and learning rate schedule (LRS) ablation.
-In PyLO, it is easy to equip VeLO and $\mu$LOM with weight decay and learning rate schedules.
-We sweep different values and find that in some cases learned optimizers improve.
-5.1 Training Vision Transformer on ImageNet-1K
-Experimental Configuration: We train ViT-Base/16 models (86M parameters) for 480 epochs (150k steps) on ImageNet-1K following established protocols (Wightman, 2019).
-We apply standard augmentation techniques (RandAugment, CutMix, Mixup) to ensure realistic training conditions that reflect practical deployment scenarios.
-We employ a batch size of 4096 and conduct these experiments on Nvidia H100 GPUs.
-Results As shown in Table 2 VeLO demonstrates competitive performance in this context, achieving 78.45% top-1 accuracy compared to AdamW’s 77.22%.
-This scenario demonstrates a case where the learned optimizer outperforms Adam.
-$\mu$LO exhibits training divergence after 65,000 steps due to its limited 5,000-step meta-training horizon, achieving only 62.14% peak validation accuracy.
-Table 2: Final Loss for Language Modeling (LM) and Final Validation Accuracy for Vision Transformer (ViT) Tasks for Medium Model Sizes Using Different Optimizers
-[…]
-$\mu$LOM 3.18 62.14
-VeLO 2.89 78.39
-AdamW + Cosine 2.91 77.22
-5.2 Language model pre-training on FineWeb
-Experimental Configuration We pre-train a $355$M parameter GPT-2 style transformers on FineWeb-EDU (Lozhkov et al., 2024) for $10$B tokens, employing a batch size of $512$ and sequences of length $1024$ (see section G for additional hyperparameter details).
-Results Table 2 reports the final loss achieved for all optimizers in our study, while figure 7 of the appendix plots the training curves.
-We observe that VeLO outperforms AdamW and $\mu$LOM, reaching the lowest final loss.
-This demonstrates that PyLO has the potential to deliver optimization performance improvements to PyTorch users.
-5.3 Combining scheduler and decoupled weight decay
-To showcase the interoperability of our library, we add learning rate scheduling and decoupled weight decay to our learned optimizer implementations in as little as $5$ lines of code.
-Although the primary objective of learned optimization is to eliminate the reliance on manually designed heuristics, we take the opportunity to investigate whether learning rate schedules and weight decay can still improve learned optimizer performance.
-Effect of Learning Rate Scheduler We equip $\mu$LOM and VeLO with a cosine annealing LRS and tune the maximum learning rate (MaxLR).
-Figure 4 reports ViT and GPT training and accuracy curves for the best-performing MaxLR while training curves for more values are reported in the appendix (F).
-We observe that $\mu$LOM benefits substantially from explicit scheduling, extending stable training from 65k to 150k steps and reaching 71% Top-1 accuracy on ImageNet and improving loss in language model pre-training.
-VeLO shows minimal scheduling improvements, suggesting different internal adaptation mechanisms.
-Effect of Weight decay We equip $\mu$LOM and VeLO with decoupled weight decay and tune the decay coefficient, $\lambda$.
-Figure 4 reports ViT and GPT training and accuracy curves for the best-performing $\lambda$ while training curves for more values are reported in the appendix (F).
-We observe that $\mu$LOM benefits from weight decay for training GPT but not VeLO.
-Neither optimizer benefits from weight decay for ViT tasks.
-6 Conclusion
-In conclusion, we have introduced PyLO, an open-source repository for state-of-the-art learned optimizers in PyTorch.
-PyLO breaks down the barriers to learned optimizers adoption by integrating directly with torch.optim.Optimizer and the HuggingFace ecosystem, allowing users to directly leverage learned optimizers in their existing code and easily download and share learned optimizer weights.
-For the best performance, PyLO implements a CUDA accelerated forward pass for the small_fc_lopt architecture, which we demonstrate has superior scaling properties and lower memory overhead than the original JAX implementation.
-When benchmarking the PyLO implementation of $\mu$LOM and VeLO against AdamW, we observed that VeLO can outperform AdamW on transformer language modeling and ViT tasks.
-We also ran ablations facilitated by PyLO’s PyTorch integration and found that learning rate schedules and weight decay can substantially improve the performance of certain learned optimizers.
-Acknowledgments
-We acknowledge support from FRQNT New Scholar [E.B.] and the FRQNT Doctoral (B2X) scholarship [B.T., A.M.]. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila.
-Appendix A Comparison to Existing Learned Optimization Frameworks
-Repository Overviews
-Open-L2O (Chen et al., 2022) offers a comprehensive benchmarking suite designed primarily for evaluating learned optimizers (L2O) across a range of optimization problems.
-It includes both model-based and model-free L2O methods, facilitating reproducibility and fair comparisons but does not focus on practical deployment for real-world tasks.
-Google’s learned_optimization (Metz et al., 2022a) is a research-oriented repository that provides a robust JAX-based framework for training, designing, evaluating, and applying learned optimizers.
-It extensively supports meta-training mechanisms, including Evolution Strategies (ES) and Predictive ES (PES), but lacks user-friendly interfaces for applying these optimizers in practical, production-oriented settings.
-PyLO (ours) addresses critical barriers that have limited the adoption of learned optimizers in the broader machine learning community by providing an accessible, performance-focused PyTorch implementation.
-PyLO features seamless integration with PyTorch workflows, CUDA acceleration for computational efficiency, and HuggingFace ecosystem integration for easy sharing of optimizer weights and community-driven development.
-Comparative Analysis
-PyLO distinguishes itself from Open-L2O and Google’s learned_optimization in several key aspects:
-  1. 1.
-    Framework Compatibility: PyLO is built entirely in PyTorch, ensuring compatibility with the predominant framework used by practitioners, unlike Google’s learned_optimization, which is JAX-based.
-  2. 2.
-    Decoupled Meta-testing: PyLO exclusively provides optimized implementations for meta-testing without the complexity and overhead associated with meta-training code, unlike the other repositories.
-  3. 3.
-    HuggingFace Integration: PyLO integrates deeply with the HuggingFace Hub, facilitating standardized sharing, versioning, and benchmarking of task-specific optimizers, a capability not provided by the other repositories.
-  4. 4.
-    Computational Efficiency: By leveraging specialized CUDA kernels, PyLO significantly reduces optimizer step overhead, enabling practical deployment for large-scale models, a substantial improvement over naive implementations.
-The strengths of PyLO position it uniquely as both researcher- and practitioner-friendly, enabling efficient exploration and practical use of state-of-the-art learned optimization techniques.
-Table 3: Comparison of Learned Optimization Repositories
-Repository HF PyTorch Decoupled Paper Github
-Open-L2O (Chen et al., 2022) Open-L2O
-Learned Optimization (Metz et al., 2022a) learned_optimization
-PyLO (ours) In submission PyLO
-Appendix B Extended challenges of L2O adoption for the Wider ML community
-This section examines the critical barriers hindering the widespread adoption of learned optimizers within the broader machine-learning community.
-Despite their demonstrated potential for reduced training wall-clock time across various domains(Metz et al., 2022a; Chen et al., 2022), several structural and technical challenges have limited the integration of learned optimizers into standard machine learning workflows.
-The main barriers can be summarized as follows:
-Several critical barriers have hindered the more widespread adoption of learned optimizers for deep learning tasks in our community: 1) overfocus on meta-learning in current frameworks, 2) absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, 4) the absence of learned optimizer benchmarking for popular machine learning tasks, and 5) the optimizer step overhead of learned optimizers.
-  1. 1.
-    The overfocus on meta-learning in current frameworks.
-  2. 2.
-    The absence of a PyTorch implementation for the most performant learned optimizers available.
-  3. 3.
-    The lack of a standardized way to effectively share learned optimizer weights.
-  4. 4.
-    The absence of learned optimizer benchmarking for popular machine learning tasks.
-  5. 5.
-    The optimizer step overhead of learned optimizers.
-Overfocus on meta-learning
-A primary obstacle to L2O adoption stems from the disparity between current learned optimization frameworks and established training paradigms.
-Existing libraries, such as the JAX (Bradbury et al., 2018)-based learned optimization framework (Metz et al., 2022a, b), prioritize meta-training capabilities over practical deployment.
-While these frameworks do provide mechanisms for utilizing pre-trained learned optimizers, the interfaces diverge significantly from those familiar to practitioners (Wightman, 2019; Wolf et al., 2020; von Platen et al., 2022).
-The evaluation and application of learned optimizers appear as secondary objectives, with insufficient emphasis on seamless integration with conventional training pipelines.
-This architectural choice highlights the necessity of decoupling meta-learning and meta-testing phases, particularly for researchers and practitioners who seek to leverage pre-trained optimization algorithms without engaging in the complexities of meta-training.
-Pytorch v.s. JAX
-A second significant barrier relates to the underlying deep learning frameworks.
-While JAX (Bradbury et al., 2018) offers powerful functional programming capabilities and high flexibility for research workflows, it lacks the mature ecosystem and widespread adoption of more established frameworks.
-In contrast, PyTorch (Paszke et al., 2019) has developed into the dominant deep learning framework with extensive community support (He, 2019) and a rich ecosystem of pre-trained models and resources (Wightman, 2019; von Platen et al., 2022; Wolf et al., 2020).
-This prevalence stems partly from PyTorch’s object-oriented architecture and modular design principles, which facilitate extension, ease-of-use, and customization.
-The absence of a comprehensive learned optimizer library that conforms to PyTorch’s optimizer interface presents a substantial obstacle, as it necessitates significant codebase modifications for implementation and experimentation.
-Consequently, there exists a pressing need for a modular L2O framework that seamlessly integrates with PyTorch’s optimizer class hierarchy.
-Such integration, particularly when combined with standardized weight serialization and distribution through established repositories, would substantially accelerate both the circulation and scaling of learned optimization techniques.
-Absence of standardized benchmarking
-A third critical limitation concerns evaluation methodologies and benchmarking procedures.
-Current L2O framework Metz et al.
-(2022a) , typically evaluate optimization algorithms across an extensive but potentially misaligned task distribution.
-These evaluations predominantly focus on training loss trajectories rather than generalization performance metrics that are of primary interest to practitioners.
-Moreover, recent works like VeLO(Metz et al., 2022b) omit crucial experimental details, including baseline hyperparameter configurations , impeding reproducibility and comparative analysis(Rezk et al., 2023).
-This deficiency underscores the necessity for standardized evaluation frameworks that enable systematic comparison of conventional and learned optimizers on practically relevant tasks.
-Such frameworks should prioritize contemporary challenges in vision and language model pre-training rather than theoretical convex optimization problems that, while analytically tractable, often fail to capture the complexities of modern neural network training dynamics.
-Sharing learned optimizer weights
-Pre-trained models for language understanding (DeepSeek-AI et al., 2025; Touvron et al., 2023), image classification (Dosovitskiy et al., 2020; Radford et al., 2021), semantic segmentation (Kirillov et al., 2023), and a number of other tasks are routinely shared via platforms such as HuggingFace Hub (HuggingFace, ), enabling practitioners to leverage computational investments made by well-resourced institutions.
-This collaborative open-model ecosystem has accelerated innovation and reduced redundancy in these domains through its expansion of access to state-of-the-art models.
-Despite this revolution in model sharing, we lack seamless sharing in learned optimizer weights.
-We suggest that foundational Learned Optimizer (LO) models can and should be integrated into the community following similar distribution mechanisms (e.g., HuggingFace) to those established for neural network weights.
-Our objective is to facilitate seamless integration of learned optimizers into existing PyTorch (Paszke et al., 2019) workflows with minimal friction, addressing the significant gap in accessibility for the predominant deep learning framework used by practitioners.
-Furthermore, by establishing infrastructure for distributing optimizer weights, we aim to catalyze a parallel ecosystem where learned optimizer rules can be shared and deployed across diverse problem domains.
-Per-step overhead of learned optimizers
-A significant obstacle to the widespread adoption of learned optimizers is the computational overhead they potentially introduce to already resource-intensive training setups.
-Deep learning workflows operate under strict time and computational constraints, with practitioners demonstrating marked reluctance to try new optimization strategies that might increase training durations, regardless of potential convergence benefits.
-This challenge necessitates meticulous implementation of learned optimization algorithms to ensure competitive computational efficiency.
-Our approach addresses this critical constraint through custom CUDA implementations that substantially reduce optimizer step time compared to native PyTorch implementations.
-By implementing the learned optimizer update as a neural network pass directly in CUDA and strategically storing intermediate activations in registers, we minimize the need to write activations to global GPU memory.
-This optimization achieves substantial reductions in wall-clock time for optimization steps and render learned optimizers practically viable within existing computational budgets, removing a significant barrier to their adoption by the broader machine-learning community.
-Appendix C Extended Code design
-We implement a modular architecture partitioned into three main components to realize our project vision.
+Interoperability Through strict adherence to the PyTorch optimizer (torch.optim.Optimizer) interface, PyLO ensures seamless integration with existing training frameworks and compatibility with popular optimizer modifiers.
+This allows practitioners to use learned optimizers with minimal modification to their existing code—including modifiers like decoupled weight decay and learning rate schedules.
+4 Library design
+We implement a modular architecture partitioned into four main components to realize our project vision.
 This design philosophy prioritizes both usability for practitioners and extensibility for researchers developing novel learned optimizers.
@@
 This can introduce substantial computational overhead for large models, creating a critical bottleneck that limits practical deployment.
-Our analysis revealed that standard PyTorch neural network modules are inadequately optimized for this specific use case.
-These modules are primarily designed for batched operations on image or language data rather than the unique computational pattern of learned optimizers, where many small MLPs must be evaluated in parallel across millions of parameters.
-Drawing inspiration from the architecture-aware optimizations (Müller et al., 2021; Dao et al., 2022) we implemented a specialized CUDA kernel for the small_fc_lopt learned optimizer.
+We emphasize that standard PyTorch neural network modules are inadequately optimized for this specific use case.
+These modules are primarily designed for batched operations on image or language data rather than the unique computational pattern of learned optimizers, where many small MLPs must be evaluated in parallel across millions or billions of parameters.
+Drawing inspiration from architecture-aware optimizations (Müller et al., 2021; Dao et al., 2022), we implemented specialized CUDA kernels for applying learned optimizers.
 This implementation strategically leverages the GPU memory hierarchy to address the primary bottleneck: memory bandwidth rather than computational throughput.
-Our CUDA implementation employs two key optimization strategies: (1) Register-Based Computation and (2) On-the-Fly Feature Computation.
+Our CUDA implementation employs two key optimization strategies: (1) Fused MLP Inference and (2) On-demand Feature Computation.
 The first utilizes GPU registers to store intermediate activations during forward propagation through the optimizer’s MLP.
@@
 The second avoids storing normalized features in global memory, recalculating them when needed, and trading redundant computation for reduced memory bandwidth.
-Appendix D Extended Experimental Results
-D.1 Benchmarking LO step times in PyLO
-Model Optimizer Batch Size Samples per sec Step time (ms) Forward time (ms) Backward time (ms) Optimizer step time(ms) Num parameters(M) Vit-B-16 AdamW 32.00 524.09 60.55 17.52 38.13 4.90 86.57 Adafactor 425.51 74.69 18.99 small_fc_lopt (naive) 39.36 812.51 756.80 small_fc_lopt (CUDA) 205.59 155.16 99.59 Vit-S-16 AdamW 32.00 1,116.12 28.17 7.65 18.27 2.25 22.05 Adafactor 711.97 44.47 18.53 small_fc_lopt (naive) 109.03 293.01 266.90 small_fc_lopt (CUDA) 382.73 83.28 57.45 Vit-L-16 AdamW 32.00 175.95 181.36 52.18 113.89 15.28 304.33 Adafactor 156.64 203.34 37.23 small_fc_lopt (naive) 11.64 2,748.52 2,582.30 small_fc_lopt (CUDA) 70.69 451.69 285.80 GPT2-125 m AdamW 4.00 17.80 224.72 73.20 144.24 7.29 125.26 Adafactor 17.02 235.07 17.76 small_fc_lopt (naive) 3.38 1,182.33 964.89 small_fc_lopt (CUDA) 11.71 341.46 124.09 GPT2-355 m AdamW 4.00 6.56 609.83 197.41 392.29 20.12 355.92 Adafactor 6.40 624.64 35.11 small_fc_lopt (naive) 1.16 3,461.96 2,872.17 small_fc_lopt (CUDA) 4.40 908.69 319.14 GPT2-1B AdamW 4.00 2.88 1,387.51 442.49 896.81 48.21 912.95 Adafactor 2.87 1,395.03 54.08 small_fc_lopt (naive) OOM OOM OOM small_fc_lopt (CUDA) 1.98 2,025.26 681.78
-Table 4: Breakdown of time spent per training step, averaged over 40 steps.
-The CUDA-based implementation reduces overhead from the learned optimizer by minimizing memory usage and optimizer time, ultimately achieving competitive throughput compared to traditional optimizers.
-We extend our analysis beyond the baseline results to examine per-step computational overhead across varying model scales representative of contemporary pretraining workloads.
-Table 4 presents a comprehensive timing breakdown across forward pass, backward pass, and optimizer step durations for both vision and language model architectures.
-Traditional optimizers including AdamW (Loshchilov & Hutter, 2019) and Adafactor (Shazeer & Stern, 2018) exhibit minimal computational overhead relative to the forward and backward passes.
-Our CUDA kernel implementations demonstrate substantial reductions in learned optimizer overhead, with improvements that scale favorably with model size.
-Notably, for ViT-L-16, our small_fc_lopt(CUDA) implementation achieves a 9× reduction in optimizer step time compared to the naive implementation.
-This performance advantage becomes increasingly pronounced at scale.
-Consistent with our earlier findings regarding memory limitations—where JAX implementations encounter out-of-memory errors for 1B parameter language models—the naive learned optimizer implementation similarly fails at this scale.
-Our CUDA implementation addresses these scalability constraints while maintaining the computational efficiency gains observed with increasing batch sizes, making learned optimization viable for large-scale pretraining scenarios.
-(a) Width
-(b) Depth
-Figure 5: Step time for deeper and wider MLPs.
-Increasing the width and depth of MLPs provides a systematic way to measure the impact of larger and more tensors on optimizer step time.
-Subfigure (a) reports optimizer step time as the hidden size of the 5-layer optimizee MLP is increased, while Subfigure (b) reports step time as the depth of a 256 hidden-layer MLP is increased.
-We observe that in each case, the CUDA implementation reduces the optimizer step time by more than an order of magnitude.
-Scaling analysis with MLP: To systematically characterize the scaling behavior of our implementation, we conduct controlled experiments using synthetic MLP architectures with varying architectural parameters.
-Figure 5 presents two complementary scaling analyses that isolate the effects of network width and depth on optimizer computational overhead.
-Figure 5(a) examines the relationship between hidden layer dimensionality and optimization time for a fixed 5-layer MLP architecture.
-The results reveal a stark performance disparity: the naive implementation of small_fc_lopt exhibits optimization times an order of magnitude greater than our CUDA implementation and encounters memory limitations beyond hidden dimensions of 256.
-While our CUDA implementation remains slower than traditional optimizers (AdamW and Adafactor), it demonstrates favorable scaling characteristics that maintain computational tractability across the tested range.
-Complementary analysis in Figure 5(b) evaluates depth scaling behavior using MLPs with fixed width of 256 units but varying layer counts (depth).
-The naive implementation consistently operates in a computationally prohibitive regime, requiring over 400ms for optimization steps, while our CUDA implementation maintains optimization times below 50ms—approaching the performance envelope of traditional optimizers despite the inherent complexity of learned optimization.
-These scaling experiments demonstrate that our CUDA implementation successfully addresses the computational bottlenecks that render naive learned optimizer implementation impractical for realistic model architectures.
+The full implementation and performance analysis are detailed in Section 5.
+Decoupled evaluation
+Providing easy-to-use examples for new open-source code can go a long way for improving adoption, as it illustrates how the library can be used and provides a starting point for exploration.
+However, it is important to decouple such examples from the learned optimizer implementation itself to keep the lines of code within PyLO to a minimum.
+As such, we provide PyLO-Examples alongside PyLO as a supporting repository for evaluating learned optimizers on popular language modelling (pre-training on FineWeb EDU) and image classification (classical ImageNet pre-training) tasks.
+Each task is implemented within a single directory to maximize readability and re-use.
+5 CUDA Implementation
+In this section, we describe the implementation details of our CUDA kernels and the motivation behind our design choices.
+We begin by analyzing the performance bottlenecks in the naive PyTorch implementation, then present our optimized approach.
+5.1 Analysis of Naive Implementation
+Figure 3: Overview of the Learned Optimizer (LO) update mechanism: The figure depicts the computation pipeline used by learned optimizers such as small_fc_lopt and VeLO.
+Model parameters ($\mathbf{\theta}$), gradients ($\mathbf{g}$), momentum (m), and second-moment accumulators (v, row & column factors of v) are combined within construct_features() to form a rich set of derivative features, including elementwise interactions such as $\mathbf{g}\times\text{momentum}$ and $\text{momentum}\times\text{factors}$ (see Appendix G for a complete description).
+These features are reduced over the parameter dimensions ($m\times n$) in compute_squared_average() to obtain normalization statistics.
+In feature_normalization(), features are normalized before being processed by the Learned Optimizer, which predicts an update direction and magnitude in apply_lo(), and this predicted update is then applied to yield the updated parameters.
+A learned optimizer step comprises several distinct operations, visualized in Figure 3.
+Consider a parameter tensor of dimensions $m\times n$.
+The naive implementation constructs a feature vector of size $mn\times d_{\text{feat}}$ where $d_{\text{feat}}$ denotes the learned optimizer’s input dimensionality.
+This construction starts after the accumulator update (update of momentum, second moment, etc) and proceeds in three stages:
+Feature Construction After updating accumulators the construct_features() procedure starts.
+The gradient $g$, parameter $p$, momentum buffer $m$, second-moment estimate $v$, and factored second-moment terms (row and column factors), are used to compute a set of derived features (see Appendix G for all features).
+Each derived feature requires individual kernel launches for element-wise operations.
+Feature normalization Before applying the learned optimizer’s neural network, features must be normalized by their squared average across all parameter elements.
+This requires: (1) computing $\mathbb{E}[\text{feature}^{2}]$ via reduction operations (compute_squared_average()), and (2) dividing each feature by $\sqrt{\mathbb{E}[\text{feature}^{2}]+\epsilon}$ (feature_normalization()).
+PyTorch’s general-purpose reduction kernels require additional kernel launches after materializing the full feature vector.
+Apply Learned Optimizer Following feature preparation and normalization, the apply_lo() procedure is invoked.
+A learned optimizer, which is implemented as a neural network, takes the normalized feature vector as input and predicts two scalar outputs: one representing the magnitude and the other the direction of the parameter update.
+These predictions are combined to compute the update according to the formula
+$\Delta\theta=\text{direction}\times e^{\text{magnitude}\times\alpha}\times\beta,$
+where $\alpha$ and $\beta$ are hyperparameters fixed at $0.01$.
+The parameter $\theta$ is then updated as $\theta\leftarrow\theta+\Delta\theta$.
+In the PyTorch-based MLP implementation, this process launches a separate kernel for each layer and stores intermediate activations in global memory.
+Consequently, memory bandwidth emerges as the primary bottleneck, resulting in underutilized GPU compute resources and reduced overall efficiency.
+The naive implementation thus requires 74-252 kernel launches per parameter tensor, with each intermediate result materialized in global memory.
+For a parameter tensor of size $mn$, this generates approximately (num_of_accum)mn(num\_of\_accum)*mn words of memory traffic and allocates $mn\times d_{\text{feat}}$ words of temporary storage, which a substantial burden for large models.
+This is evidenced by the profiler visualization as shown in Figure 4.
+Figure 4: Comparison of CUDA kernel execution characteristics for the learned optimizers small_fc_lopt and VeLO, shown for both naïve and fused implementations applied to a single layer MLP (shape [1000x1000]) with no bias.
+(Left) The total cumulative kernel execution time (blue) and the number of kernel launches (orange) reveal that the naïve implementations spend substantially more time executing a far greater number of small kernels, due to performing each arithmetic or reduction step as an independent operation.
+In contrast, the fused CUDA versions combine these operations into fewer, more computationally dense kernels, reducing both launch overhead and memory bandwidth pressure.
+(Right) Breakdown of cumulative kernel time by kernel type for each optimizer shows that the fused versions (bolded) consolidate many repetitive operations into specialized fused kernels.
+This pattern is consistent across both VeLO and small_fc_lopt, demonstrating that kernel fusion is broadly beneficial for improving learned optimizer efficiency.
+5.2 Optimized Kernel Design
+Our optimization strategy rests on two key observations: (1) feature construction can be performed in a single pass over the parameter tensor, and (2) the neural network forward pass can be fused with feature construction to eliminate intermediate storage.
+To this end, we implement a two-pass method in which we decompose the learned optimizer step into two kernels:
+Kernel 1: Feature statistics collection.
+We perform a single pass over all parameters, computing features on-the-fly in registers and accumulating their squared values for normalization.
+Rather than materializing the full feature tensor, we maintain only a $d_{\text{feat}}$-dimensional accumulator per thread.
+Reduction proceeds hierarchically: thread-local accumulators are reduced within each warp using shuffle instructions, warp results are aggregated via shared memory, and block-level results are combined through atomic operations to global memory.
+This approach computes the required statistics using only $O(d_{\text{feat}})$ shared memory per block and $O(d_{\text{feat}})$ global memory total, independent of parameter count.
+This combines the construct_features() and compute_squared_average() procedures.
+Pseudocode for Kernel 1
+// Phase 1: Fused Compute Statistics
+Kernel_1(g, p, m, v, ...) {
+thread_accum[d_feat] = {0}; // Register array
+for (i in grid_stride_loop) {
+features = compute_features(g[i], p[i], m[i], v[i], ...);
+thread_accum += features * features;
+warp_reduce(thread_accum); // Shuffle operations
+block_reduce(warp_results); // Shared memory
+atomic_add(global_stats, block_result);
+Kernel 2: Second Fused feature construction and network application.
+With normalized statistics available in global memory, we perform a second pass that: (1) loads the normalization factors into shared memory for efficient broadcast, (2) constructs features in registers from parameter and accumulator values for the second time, (3) evaluates the learned optimizer network in-line, and (4) applies the computed update.
+No intermediate tensors are materialized—all computation flows through the register file.
+This combines construct_features(),feature_normalization() and apply_lo() procedures.
+Memory hierarchy exploitation Our design carefully manages data placement across the GPU memory hierarchy:
+Registers: Feature vectors ($d_{\text{feat}}\approx 39$ elements) and network activations reside entirely in registers during computation, providing single-cycle access latency.
+Shared memory: The $d_{\text{feat}}$-dimensional normalization statistics are loaded once per thread block and broadcast to all threads, amortizing the cost of global memory access across all threads in the block.
+Global memory: Only parameter values, gradients, and optimizer state buffers are read from global memory, and only updated parameters are written back.
+This reduces memory traffic substantially per optimization step.
+This design eliminates kernel launch overhead (reducing from 74-252 to 30-114 kernel invocations, see Figure 4), removes intermediate allocations (reducing memory footprint by $mn\times d_{\text{feat}}$ words), and minimizes memory bandwidth consumption (Data movement is minimized as only accumulators are read and final outputs are written).
+The result is an implementation that achieves good memory bandwidth utilization while maintaining numerical equivalence to the naive approach.
+Pseudocode for Kernel 2
+// Phase 2: Fused Apply
+Kernel_2(g, p, m, v, ..., global_stats) {
+__shared__ normalized_stats[d_feat];
+if (tid < d_feat) {
+normalized_stats[tid] = rsqrt(global_stats[tid] / N); // N = number of params
+__syncthreads();
+for (i in grid_stride_loop) {
+// All computation in registers
+features[d_feat] = compute_features(g[i], p[i], m[i], v[i], ...);
+features *= normalized_stats; // Broadcast from SMEM
+// Inline MLP (weights in __ldg cache)
+hidden = relu(Linear1(features));
+hidden = relu(Linear2(hidden));
+direction, magnitude = Linear3(hidden);
+update = direction * exp(magnitude * alpha) * beta;
+p[i] -= update;
+6 Benchmarking LO step times in PyLO
+In this section, we establish the per-step overhead of learned optimizers for common pre-training workloads and showcase the improved scaling behavior of PyLO’s CUDA-accelerated learned optimizer steps.
+Our goal is to report learned optimizer overhead relative to hand-designed optimizers for practical tasks and to illustrate how the overhead scales as the model size is increased.
+To this end, we benchmark and record the forward, backward, and optimizer step times for training vision transformers and transformer language models of different sizes on a single 8080GB A100 GPU.
+Learned Optimizer overhead shrinks as batch size increases.
+Figure 1 reports timings at different batch sizes for ViT-B/16.
+We observe that as batch size is increased, the overhead of the learned optimizer’s step relative to the forward and backward passes diminishes.
+This indicates that large-scale training with prolonged per-step compute may be particularly well-suited for learned optimizers, as the optimizer cost can be effectively amortized.
+This trend is further evidenced in Table 2 for both vision and language modeling, which presents timing comparisons across optimizers for ViT-B/16 and a medium-sized GPT-2 model (355M parameters).
+We find that relative throughput compared to Adam improves with model scale, holding true for language modeling as well.
+Additionally, per-step time for small_fc_lopt (CUDA) scales with model size, yet our CUDA kernels deliver substantial speedups over the naive baseline.
+For ViT-B/16, small_fc_lopt (CUDA) achieves an 86% reduction in optimizer step time versus the naive version; for the GPT-2 task, the reduction is 88%.
+(Full results, including larger ViT and GPT variants, are provided in Appendix F.)
+Figure 5: Step Time Scaling of Learned Optimizers.
+We present a comparison of optimizer step time between our custom CUDA implementation and the original JAX versions of small_fc_lopt and VeLO, evaluated during the training of GPT-2 style transformer models across a range of model sizes.
+The results show that our CUDA implementation not only achieves substantially lower step times but also maintains this advantage as model size increases, enabling more efficient scaling to larger architectures.
+Model Optimizer BS/SL Fwd (ms) Bwd (ms) Opt step (ms) #Params (M)
+ViT-B/16 AdamLoshchilov & Hutter (2019) 32/197 17.52 38.13 4.90 86.57
+AdafactorShazeer & Stern (2018) 18.99
+small_fc_lopt (naive) 756.80
+small_fc_lopt (CUDA) 99.59 (-86%)
+VeLO (naive) 585.11
+VeLO (CUDA) 113.58 (-80%)
+GPT2-355M AdamLoshchilov & Hutter (2019) 4/1024 197.41 392.29 20.12 355.92
+AdafactorShazeer & Stern (2018) 35.11
+small_fc_lopt (naive) 2872.17
+small_fc_lopt (CUDA) 319.14 (-88%)
+VeLO (naive) 2378.93
+VeLO (CUDA) 284.37 (-88%)
+Table 2: Timing Results Across Optimizers for ViT and GPT Models.
+We report optimizer step times for Vision Transformer (ViT-B/16) and a GPT-2 style model with 355M parameters, using realistic batch sizes (BS) and sequence lengths (SL) that fit within the memory constraints of a single A100 GPU.
+The results demonstrate that our custom CUDA kernel significantly reduce optimizer step time (reduction shown in red) compared to naive implementation, enabling more efficient training in practical settings.
+(a) 1-layer MLP
+(b) 256-width MLP
+(c) 4096-width MLP
+Figure 6: Scaling behavior of CUDA implementation with varying MLP configurations We evaluate optimizer step time across various MLPs with controlled width and depth.
+Subfigure (a) reports optimizer step time as the hidden size of the 1-layer optimizee MLP is increased.
+Subfigure (b) reports step time as the depth of a 256 hidden-dimension optimizee MLP is increased.
+Subfigure (c) reports step time as the depth of a $4096$ hidden-dimension optimizee is increased.
+Missing datapoints correspond to OOM errors.
+We observe that in each case, our CUDA implementation reduces the optimizer step time by an order of magnitude relative to the naive implementation.
+Scaling behavior of our CUDA implementation vs the JAX baseline Figure 5 reports the scaling behavior w.r.t.
+transformer size when trained with different implementations of small_fc_lopt (we use hidden size $32$ as in Thérien et al.
+(2024)) and VeLOMetz et al.
+(2022b).
+JAX corresponds to the original implementation of Metz et al.
+(2022a), which uses the JAX compiler, while CUDA is the fastest implementation that we make available in PyLO.
+We observe that our CUDA implementations results in a substantial memory savings as it can optimize a 1B parameter transformer, while the JAX implementation runs out of memory for an 80 GB A100 GPU.
+Moreover, in cases where the JAX implementation does not run out of memory, it is outperformed by our CUDA implementation, resulting in a $2\times$ reduction in step time.
+Isolating Scaling behavior of our CUDA implementation with MLP taks To systematically characterize the scaling behavior of our implementation, we conduct controlled experiments using synthetic MLP architectures with varying numbers of layers and hidden dimensions.
+This allows us to explicitly control for tensor sizes and the number of tensors to measure the effect on our implementation.
+Figure 6 presents three complementary scaling analyses that isolate the effects of network width and depth on optimizer computational overhead.
+Figure 6(a) examines the relationship between hidden layer dimensionality and optimization time for a fixed 1-layer MLP architecture.
+The results reveal a stark performance disparity: the naive implementations of learned optimizers exhibit optimization times an order of magnitude greater than our CUDA implementation.
+While our CUDA implementations remains slower than traditional optimizers, it demonstrates favorable scaling characteristics.
+We further enhance this on H100 hardware in Appendix A.
+Complementary analysis in Figure 6(b),(c) evaluates depth scaling behavior using MLPs with fixed width of 256 and 4096 units but varying layer counts (depth).
+The naive implementation consistently operates in a computationally prohibitive regime, requiring over 400ms for optimization steps, while our CUDA implementation maintains optimization times approaching the performance envelope of traditional optimizers despite the inherent complexity of learned optimization.
+These scaling experiments demonstrate that our CUDA implementation successfully addresses the computational bottlenecks that render naive learned optimizer implementations impractical for realistic model architectures.
+7 Distributed Optimizer Step
+It is possible to further reduce optimizer step time by distributing computation across multiple devices.
+So far, our analysis has focused on optimizer step times for fully replicated data-parallel training where the gradient is first all-reduced and, subsequently, the optimizer step for the entire model is performed redundantly and simultaneously on all devices.
+While not ideal, this approach is reasonable for coordinate-wise optimizers like Adam, which require little computation during the optimizer step.
+However, optimizers like ShampooGupta et al.
+(2018), MuonJordan et al.
+(2024), or Learned optimizers Metz et al.
+(2022a; b), that perform significant computation at each step, can benefit from distributing the optimizer step across multiple devices.
+As illustrated in figure 7, the optimizer step can effectively be distributed across devices by splitting the all-reduce operation into its parts: (1) reduce scatter the gradients evenly for each parameter tensor across devices, (2) perform the optimizer step non-redundantly across all devices, and (3) all-gather the updated model parameters.
+Table 3: Comparing improvements from distributed optimizer steps.
+We report the time taken for the forward and backward pass, communication, and optimizer step for different optimizers when training a 125M transformer language model across four H100 GPUs.
+All times are reported in milliseconds (ms).
+The all-reduce column performs an all-reduce of the gradient before the optimizer step, while reduce-scatter distributes the optimizer step across devices (Figure 7 illustrates the difference between these steps.).
+We observe that using a distributed optimizer step can reduce learned optimizer step times by up to 31.86 ms.
+Optimizer All-reduce (ms) Reduce-scatter (ms) Difference
+small_fc_lopt CUDA 136.53 104.67 31.86
+VeLO CUDA 127.67 108.71 18.96
+Muon 106.92 98.10 8.82
+Adam 99.93 95.93 4.00
+Improvements from using a distributed optimizer step Table 3 reports the time taken for the forward and backward pass, communication, and optimizer step for different optimizers when training a 125M transformer language model across four H100 GPUs.
+We observe that distributing the optimizer step can lead to substantial improvements for all optimizers, but that learned optimizers improve the most from a distributed step: by $31.86$ ms for small_fc_lopt and $18.96$ ms for VeLO.
+This reduces the overhead relative to Adam for these optimizers to just $9\%$ and $13\%$, respectively.
+These results suggest that, in large batch training settings, the overhead or learned optimizers relative to Adam and Muon can further be reduced by distributing the optimizer step over more devices.
+Sharded Optimizer States.
+Our analysis shows that tensor-level reduce-scatter significantly accelerates distributed optimizer steps.
+Further gains in memory and compute are possible by sharding optimizer states across devices, as in ZeRO Rajbhandari et al.
+(2020) (i.e., fully sharded data parallelism Zhao et al.
+(2023)).
+However, completing an optimizer step may require reconstructing full accumulator tensors via all-to-all and all-gather operations.
+Since learned optimizers currently operate per-parameter (except normalization), the costly all-to-all can be replaced with a distributed normalization step that communicates only $d_{\text{feat}}$ scalars.
+We leave sharded learned optimizers for future work.
+8 Demonstrating the real-world use of PyLO
+[…]
+Figure 7: Overlapping Optimizer steps with Communication and distributed Optimizer steps.
+[…]
+(b) LM VeLO
+[…]
+(d) ViT VeLO
+[…]
+8.1 Training Vision Transformer on ImageNet-1K
+Experimental Configuration: We train ViT-Base/16 models (86M parameters) for 480 epochs (150k steps) on ImageNet-1K following established protocols Wightman (2019).
+We apply standard augmentation techniques (RandAugment, CutMix, Mixup) to ensure realistic training conditions that reflect practical deployment scenarios.
+We employ a batch size of 4096 and conduct these experiments on Nvidia H100 GPUs.
+[…]
+Table 4: Final Loss for Language Modeling (LM) and Final Validation Accuracy for Vision Transformer (ViT) Tasks for Medium Model Sizes Using Different Optimizers
+Optimizer LM Loss $\downarrow\;$(355M) ViT Acc. $\uparrow\;$(86M)
+[…]
+VeLO 2.89 78.39
+Adam + Cosine 2.91 77.22
+8.2 Language model pre-training on FineWeb
+Experimental Configuration
+We pre-train a 355355M parameter GPT-2 style transformers on FineWeb-EDU Lozhkov et al.
+(2024) for 1010B tokens, employing a batch size of $512$ and sequences of length $1024$ (see section D for additional hyperparameter details).r more details.
+Results
+[…]
+8.3 Combining scheduler and decoupled weight decay
+To showcase the interoperability of our library, we add learning rate scheduling and decoupled weight decay to our learned optimizer implementations in as little as $5$ lines of code.
+Although the primary objective of learned optimization is to eliminate the reliance on manually designed heuristics, we take the opportunity to investigate whether learning rate schedules and weight decay can still improve learned optimizer performance.
+Effect of Learning Rate Scheduler
+[…]
+Effect of Weight decay
+[…]
+9 Conclusion
+We have presented PyLO, an open-source PyTorch library that removes key barriers to the practical use of learned optimizers in large-scale deep learning.
+PyLO offers drop-in implementations of state-of-the-art methods such as VeLO and small_fc_lopt through the standard torch.optim.Optimizer interface, with seamless Hugging Face Hub integration for model sharing.
+Central to our systems contributions are custom CUDA kernels that fuse feature construction, normalization, and MLP inference, and a distributed implementation that distributes tensors across multiple devices.
+This achieves 86–88% lower overhead than naive PyTorch, and can efficiently scale to training a 1B-parameter transformer on a single A100 GPU.
+Our evaluation shows that PyLO reduces learned optimizer overhead to acceptable levels at scale and they can deliver competitive accuracy on ImageNet and large-scale language pretraining.
+As future work, we envision PyLO enabling (1) hardware-optimizer co-design (2) behavior-cloned distillation of expensive second-order optimizers such as SOAPVyas et al.
+(2025).
+PyLO thus establishes a robust foundation for community-driven research and practical deployment of learned optimizers in modern ML systems.
+Appendix A Enhanced Performance on NVIDIA H100 GPUs
+To further demonstrate the scalability of our approach on state-of-the-art hardware, we evaluate its performance on NVIDIA H100 GPUs, which offer advanced features.
+These accelerators are particularly beneficial for large matrix operations, such as those with dimensions $8192\times 8192$—corresponding to the hidden size of a 70B-parameter transformer model.
+In our enhanced implementation (denoted small_fc_lopt-PyLO-CUDA++), we introduce the following optimizations:
+  • Reduction in the number of global MLP reads to minimize HBM accesses;
+  • Asynchronous memory copies to overlap data movement with computation; and
+  • Full utilization of tensor cores and WGMMA instructions for high-throughput matrix multiplications on large tensors.
+These modifications substantially reduce the optimizer step time, effectively closing the performance gap relative to state-of-the-art baselines for large-scale workloads.
+The results are summarized in Figure 9.
+Figure 9: Optimizer step time comparison on NVIDIA H100 GPUs for $8192\times 8192$ matrices.
+Our enhanced PyLO-CUDA++ implementation (dark green) achieves a $2.1\times$ speedup over the baseline PyLO-CUDA (light green).
+Muon (pink) and Adam (orange) given for reference
+Appendix B Training Curves for the Experiments
 (a) ViT small patch 16
@@
 (c) ViT large patch 16
-Figure 6: Vision Transformer training dynamics.
-Training loss curves for Vision Transformers trained on ImageNet-1k classification across three model configurations: (a) ViT-small with patch size 16, (b) ViT-base with patch size 16, and (c) ViT-large with patch size 16.
-We compare AdamW with cosine learning rate scheduling (blue), VeLO learned optimizer (orange), and $\mu$LO learned optimizer (black) over 150,000 training steps with batch size 4,096.
-$\mu$LO exhibits early training instability across all scales, while VeLO demonstrates competitive or superior convergence properties, particularly evident in the base model configuration where it achieves lower final loss than the AdamW baseline.
-Appendix E Training Curves for the Experiments
-This section presents a comprehensive empirical evaluation comparing learned optimizers VeLO (Metz et al., 2022b) and $\mu$LO (Thérien et al., 2024) against the standard AdamW optimizer (Loshchilov & Hutter, 2019; Kingma & Ba, 2017) with cosine learning rate scheduling (Loshchilov & Hutter, 2017).
-Our evaluation spans Vision Transformer architectures of varying scales and GPT-2 models, extending beyond the main paper’s validation accuracy results for ViT-Base/16 and final loss metrics for GPT-2 to provide detailed training dynamics across multiple configurations.
-Vision Transformer Experiments: We evaluate three ViT configurations—Small, Base, and Large with patch size 16—on ImageNet-1k classification tasks.
-All configurations maintain consistent training protocols with 150,000 optimization steps and batch sizes of 4,096 across model variants, using hyperparameters detailed in Table 6.
-The extended evaluation presented in Figure 6 reveals critical scale-dependent optimization characteristics.
-MuLO demonstrates pronounced early divergence patterns across both small and large model configurations, reinforcing our hypothesis that its constrained meta-training horizon limits generalization to extended optimization trajectories.
-In contrast, VeLO exhibits remarkable consistency in convergence properties across the full spectrum of model scales examined, maintaining stable training dynamics while achieving competitive performance metrics relative to the AdamW baseline.
-Language Modeling Experiments For GPT-2 pretraining on FineWeb-EDU data, we train models for 19,073 steps with batch size 512 and sequence length 1024.
-Complete hyperparameters are presented in Table 8, with full training curves shown in Figure 7.
-VeLO demonstrates competitive performance against AdamW with cosine scheduling, while MuLO exhibits significantly degraded performance across the extended training horizon, consistent with the stability patterns observed in vision tasks.
+[…]
 (a) 355M LM
@@
 We train all models for 10B tokens of FineWeb-EDU data.
-Appendix F Extended Ablation Studies
+[…]
+Vision Transformer Experiments: We evaluate three ViT configurations—Small, Base, and Large with patch size 16—on ImageNet-1k classification tasks.
+All configurations maintain consistent training protocols with 150,000 optimization steps and batch sizes of 4,096 across model variants, using hyperparameters detailed in Table 6.
+The extended evaluation presented in Figure 10 reveals critical scale-dependent optimization characteristics.
+MuLO demonstrates pronounced early divergence patterns across both small and large model configurations, reinforcing our hypothesis that its constrained meta-training horizon limits generalization to extended optimization trajectories.
+In contrast, VeLO exhibits remarkable consistency in convergence properties across the full spectrum of model scales examined, maintaining stable training dynamics while achieving competitive performance metrics relative to the Adam baseline.
+Language Modeling Experiments For GPT-2 pretraining on FineWeb-EDU data, we train models for 19,073 steps with batch size 512 and sequence length 1024.
+Complete hyperparameters are presented in Table 8, with full training curves shown in Figure 11.
+VeLO demonstrates competitive performance against Adam with cosine scheduling, while MuLO exhibits significantly degraded performance across the extended training horizon, consistent with the stability patterns observed in vision tasks.
+Appendix C Extended Ablation Studies
 This section presents comprehensive training curves across all hyperparameter configurations examined in our weight decay and cosine annealing ablations.
 Our analysis encompasses systematic sweeps of weight decay coefficients and maximum learning rates for cosine annealing schedules.
-Vision Transformer Ablations Figures 8 and 9 demonstrate the variation of training dynamics with varying maximum learning rates and weight decay values for the ViT-B/16 model.
+Vision Transformer Ablations Figures 12 and 13 demonstrate the variation of training dynamics with varying maximum learning rates and weight decay values for the ViT-B/16 model.
 The results reveal that VeLO exhibits minimal sensitivity to weight decay and scheduler modifications in this experimental setup, suggesting robust inherent regularization properties.
@@
 Language Modeling Ablations
-The language modeling experiments, presented in Figures 10 and 11, demonstrate more pronounced sensitivity to weight decay and scheduling compared to vision tasks.
+The language modeling experiments, presented in Figures 14 and 15, demonstrate more pronounced sensitivity to weight decay and scheduling compared to vision tasks.
 VeLO shows measurable performance improvements when augmented with weight decay and scheduling for specific parameter configurations.
@@
 Architectural Configuration Effects
-Figure 12 presents an additional ablation examining the impact of separate versus concatenated query-key-value (QKV) matrices on the optimizee’s training loss.
-We observe that separating the QKV matrices in attention blocks improves performance across VeLO and $\mu$LOM.
+[…]
 (a) VeLO Decoupled Weight decay ablation
@@
 Figure 12: Training ViT-B-16 with VeLO on Imagenet 1k with decoupled weight decay and cosine annealing. We see no improvement in using weight decay or Scheduler for VeLO in this setup
-(a) $\mu$LO Decoupled Weight decay ablation
-(b) $\mu$LO Cosine annealing Ablation
-Figure 9: Training ViT-B/16 with $\mu$LO on Imagenet1k with decoupled weight decay and cosine annealing.
-We see that the training horizon of $\mu$LO is improved with using a scheduler that helps it to run up to 150000 training steps.
-We also see that weight decay did not have substantial improvement
+[…]
+[…]
+[…]
 (a) VeLO Decoupled Weight decay ablation
@@
 We train all models for $10B$ tokens of FineWeb-EDU data.
-(a) $\mu$LO Decoupled Weight decay ablation
-(b) $\mu$LO Cosine annealing Ablation
-(c) $\mu$LO LR Tuning Ablation
+[…]
+[…]
+[…]
 Figure 15: GPT2 Pre-training VeLO ablations: decoupled weight decay, LR-tuning, and cosine annealing.
@@
 (a) VeLO
-(b) $\mu$LO
-Figure 12: Separate QKV weight ablation.
-We ablate using concatenated or separate QKV matrices in attention layers.
-From the perspective of the learned optimizer, these two settings will be treated differently.
-We observe that performance improves for both $\mu$LOM and VeLO when QKV matrices are separated.
-Appendix G Hyperparameters for Experiments
+[…]
+[…]
+Appendix D Hyperparameters for Experiments
 Table 5: Vision Transformer Training Hyperparameters.
@@
 Training Epochs $480$
-Optimizer (for baseline) AdamW
+Optimizer (for baseline) Adam
 Learning Rate ($\eta$) $4\times 10^{-3}$
@@
 ViT-Small/16
-Parameters $22.05$M
+Parameters 22.0522.05M
 Hidden Dimension ($d_{\text{model}}$) $384$
 MLP Hidden Dimension $1536$
-Number of Heads $6$
+Number of Heads 66
 Number of Layers $12$
@@
 ViT-Base/16
-Parameters $86.57$M
+Parameters 86.5786.57M
 Hidden Dimension ($d_{\text{model}}$) $768$
@@
 ViT-Large/16
-Parameters $307.40$M
+Parameters 307.40307.40M
 Hidden Dimension ($d_{\text{model}}$) $1024$
@@
 Training Iterations $19073$
-Optimizer (for baseline) AdamW
+Optimizer (for baseline) Adam
 Learning Rate Scheduler (for baseline) Cosine
@@
 FFN Hidden Dimension $8192$
+Appendix E Validation of PyLO Implementation Against Original JAX Codebase
+[…]
+The results demonstrate strong agreement between the two implementations across both optimizers.
+As shown in Figure 17, the training curves exhibit nearly identical convergence patterns across all 5 different runs.
+Minor variations arise from inherent differences in how PyTorch and JAX compute gradients.
+Accounting for this, we confirm that our PyTorch implementation faithfully reproduces the behavior of the original JAX code.
+We further plan to compare both implementations across a range of models and tasks.
+[…]
+(b) VeLO implementation comparison
+[…]
+Appendix F Extended per time step results
+Model Optimizer Batch Size Samples/sec Step time (ms) Fwd (ms) Bwd (ms) Opt step (ms) Params (M) ViT-B-16 AdamW 32 524.09 60.55 17.52 38.13 4.90 86.57 Adafactor 425.51 74.69 18.99 small_fc_lopt (naive) 39.36 812.51 756.80 small_fc_lopt (CUDA) 205.59 155.16 99.59 VeLO (naive) 49.73 644.45 588.80 VeLO (CUDA) 191.18 168.61 112.97 ViT-S-16 AdamW 32 1,116.12 28.17 7.65 18.27 2.25 22.05 Adafactor 711.97 44.47 18.53 small_fc_lopt (naive) 109.03 293.01 266.90 small_fc_lopt (CUDA) 382.73 83.28 57.45 VeLO (naive) 94.84 337.68 311.77 VeLO (CUDA) 242.96 132.49 106.57 ViT-L-16 AdamW 32 175.95 181.36 52.18 113.89 15.28 304.33 Adafactor 156.64 203.34 37.23 small_fc_lopt (naive) 11.64 2,748.52 2,582.30 small_fc_lopt (CUDA) 70.69 451.69 285.80 VeLO (naive) 15.41 2,077.28 1,911.20 VeLO (CUDA) 76.74 418.29 252.21 GPT2-125 M AdamW 4 17.80 224.72 73.20 144.24 7.29 125.26 Adafactor 17.02 235.07 17.76 small_fc_lopt (naive) 3.38 1,182.33 964.89 small_fc_lopt (CUDA) 11.71 341.46 124.09 VeLO (naive) 3.48 1,148.81 931.37 VeLO (CUDA) 11.25 355.49 138.05 GPT2-410 M AdamW 4 6.56 609.83 197.41 392.29 20.12 355.92 Adafactor 6.40 624.64 35.11 small_fc_lopt (naive) 1.16 3,461.96 2,872.17 small_fc_lopt (CUDA) 4.40 908.69 319.14 VeLO (naive) 1.34 2,976.82 2,387.12 VeLO (CUDA) 4.58 873.41 283.71 GPT2-1B AdamW 4 2.88 1,387.51 442.49 896.81 48.21 912.95 Adafactor 2.87 1,395.03 54.08 small_fc_lopt (naive) OOM OOM OOM small_fc_lopt (CUDA) 1.98 2,025.26 681.78 VeLO (naive) OOM OOM OOM VeLO (CUDA) 2.24 1,782.43 443.13
+Table 9: Per-step training breakdown (averaged over 40 steps).
+We compare AdamW, Adafactor, small_fc_lopt, and the VeLO optimizer in both naive (PyTorch) and CUDA-accelerated implementations.
+CUDA versions significantly reduce optimizer step time and memory overhead, enabling competitive throughput—especially for large models.
+Appendix G Features for learned optimizer input
+In the following section, we introduce the input features for small_fc_lopt and VeLO in Tables 10 and 11, respectively.
+Table 10: MLP input features for small_fc_lopt.
+We replicated the table from Thérien et al.
+(2024) for the reader’s convenience with only slight modifications; We also keep the same figure caption.
+All the coefficients, $\beta_{i}$, are learnable parameters adjusted during meta-optimization.
+All feature calculations and scalings are reported for a hidden weight matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$ in an optimizee network following our proposed $\mu$-parameterization.
+Here, $n$ is the width and $m=kn$ for some constant $k\in\mathbb{R}$.
+Notation.
+The table will use $\nabla_{t,i}$ or $\nabla_{t,j}$ to indicate the variable’s dependence on time $t$ and coefficient $\beta_{i}$ or $\beta_{j}$, respectively.
+$(\nabla_{t,j})_{r,c}$ will designate indexing into row $r$ and column $c$ of the quantity $\nabla_{t,j}$.
+Type # Description Accumulator Update/Equation
+Accumulators 3 Momentum accumulators with coefficients $\beta_{i},i\in\{1,2,3\}$. ${\bm{M}}_{t,i}=\beta_{i}{\bm{M}}_{t-1,i}+(1-\beta_{i})\nabla_{t}$
+1 Second moment accumulator with coefficient $\beta_{4}$. ${\bm{V}}_{t}=\beta_{4}{\bm{V}}_{t-1}+(1-\beta_{4})\nabla_{t}^{2}$
+3 Adafactor row accumulator with coefficients $\beta_{i},i\in\{5,6,7\}$. ${\bm{r}}_{t,i}=\beta_{i}{\bm{r}}_{t-1,i}+(1-\beta_{i})~\texttt{row\_mean}(\nabla_{t}^{2})$
+3 Adafactor accumulator with coefficients $\beta_{i},i\in\{5,6,7\}$. ${\bm{c}}_{t,i}=\beta_{i}{\bm{c}}_{t-1,i}+(1-\beta_{i})~\texttt{col\_mean}(\nabla_{t}^{2})$
+Accumulator Features 3 Momentum values normalized by the square root of the second moment for $i\in\{5,6,7\}$. $\frac{{\bm{M}}_{t,i}}{\sqrt{{\bm{V}}_{t}}}$
+1 The reciprocal square root of the second moment value. $\frac{1}{\sqrt{{\bm{V}}}}$
+6 The reciprocal square root of the Adafactor accumulators. $\frac{1}{\sqrt{{\bm{r}}_{t,i}}}\textsc{ or }\frac{1}{\sqrt{{\bm{c}}_{t,i}}}$
+3 Adafactor gradient features for $i\in\{5,6,7\}$. $\nabla_{t}\cdot\sqrt{\frac{\frac{1}{m}\sum_{h=1}^{m}({\bm{r}}_{t,i})_{h}}{{\bm{r}}_{t,i}{\bm{c}}_{t,i}^{T}}}$
+3 Adafactor momentum features for $i,j\in\{(5,1),(6,2),(7,3)\}$. ${\bm{M}}_{t,j}\cdot\sqrt{\frac{\frac{1}{m}\sum_{h=1}^{m}({\bm{r}}_{t,i})_{h}}{{\bm{r}}_{t,i}{\bm{c}}_{t,i}^{T}}}$
+Time Features 11 Time features for $x\in\{1,3,10,30,100,300,1000,3000,10^{4},3\cdot 10^{4},10^{5}\}$. $\tanh{\left(\frac{t}{x}\right)}$
+Parameters 1 Parameter value. ${\bm{W}}_{t}$
+1 Gradient value. $\nabla_{t}$
+Total $39$
+Table 11: MLP input features for VeLO.
+We replicated the table from Thérien et al.
+(2024) for the reader’s convenience with only slight modifications; We also keep the same figure caption.
+All feature calculations and scalings are reported for a hidden weight matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$ in an optimizee network following our proposed $\mu$-parameterization.
+Here, $n$ is the width and $m=kn$ for some constant $k\in\mathbb{R}$.
+Notation.
+The table will use $\nabla_{t,i}$ or $\nabla_{t,j}$ to indicate the variable’s dependence on time $t$ and coefficient $\beta_{i}$ or $\beta_{j}$, respectively.
+$(\nabla_{t,j})_{r,c}$ will designate indexing into row $r$ and column $c$ of the quantity $\nabla_{t,j}$.
+Type # Description Accumulator Update/Equation
+Accumulators 3 Momentum accumulators with coefficients $\beta_{i},i\in\{1,2,3\}$. ${\bm{M}}_{t,i}=\beta_{i}{\bm{M}}_{t-1,i}+(1-\beta_{i})\nabla_{t}$
+1 Second moment accumulator with coefficient $\beta_{4}$. ${\bm{V}}_{t}=\beta_{4}{\bm{V}}_{t-1}+(1-\beta_{4})\nabla_{t}^{2}$
+3 Adafactor row accumulator with coefficients $\beta_{i},i\in\{5,6,7\}$. ${\bm{r}}_{t,i}=\beta_{i}{\bm{r}}_{t-1,i}+(1-\beta_{i})~\texttt{row\_mean}(\nabla_{t}^{2})$
+3 Adafactor accumulator with coefficients $\beta_{i},i\in\{5,6,7\}$. ${\bm{c}}_{t,i}=\beta_{i}{\bm{c}}_{t-1,i}+(1-\beta_{i})~\texttt{col\_mean}(\nabla_{t}^{2})$
+Accumulator Features 3 Momentum values normalized by the square root of the second moment for $i\in\{5,6,7\}$. $\frac{{\bm{M}}_{t,i}}{\sqrt{{\bm{V}}_{t}}}$
+1 The reciprocal square root of the second moment value. $\frac{1}{\sqrt{{\bm{V}}}}$
+6 The reciprocal square root of the Adafactor accumulators. $\frac{1}{\sqrt{{\bm{r}}_{t,i}}}\textsc{ or }\frac{1}{\sqrt{{\bm{c}}_{t,i}}}$
+3 Adafactor gradient features for $i\in\{5,6,7\}$. $\nabla_{t}\cdot\sqrt{\frac{\frac{1}{m}\sum_{h=1}^{m}({\bm{r}}_{t,i})_{h}}{{\bm{r}}_{t,i}{\bm{c}}_{t,i}^{T}}}$
+3 Adafactor momentum features for $i,j\in\{(5,1),(6,2),(7,3)\}$. ${\bm{M}}_{t,j}\cdot\sqrt{\frac{\frac{1}{m}\sum_{h=1}^{m}({\bm{r}}_{t,i})_{h}}{{\bm{r}}_{t,i}{\bm{c}}_{t,i}^{T}}}$
+Parameters 1 Parameter value. ${\bm{W}}_{t}$
+1 Gradient value. $\nabla_{t}$
+1 Clipped gradient value. $\texttt{clip}(\nabla_{t},-0.1,0.1)$
+Total $29$